### v1

# DoubleAgents: Exploring Mechanisms of Building Trust with Proactive AI

Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people actually trust systems that act on their behalf. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their reliance before live use. We evaluate DoubleAgents in a two-day lab study (n=10), two deployments (n=2), and a technical evaluation. Results show that participants initially hesitated to delegate but grew more reliant as they experienced transparency, control, and adaptive learning during simulated cases. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that the effort required scaled appropriately with task complexity and contextual data. We contribute trust-by-design patterns and mechanisms for proactive AI—consistency, controllability, and explainability—along with simulation as a safe path to build and calibrate trust over time.
generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
copyright: acmcopyrightjournalyear: 2025conference: ; ; booktitle: Proceedings of the 2025price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
1. Introduction
Coordinating human effort is essential for tasks at every scale. Event planning involves finding suitable times, locations, vendors, and transportation. Warehouses must assign workers with the right skills and experience to shifts they are available for. Even organizing a speaker series requires emailing potential speakers, collecting their availability, and assigning them to appropriate slots. These problems involve both discovering people’s availability and solving the constraints of assigning individuals to limited roles. Human coordinators face time-consuming and intricate communication challenges: people may hesitate to disclose their full availability, delay responses, or provide only tentative answers such as “I might be able to come Saturday, but I’m not sure yet.” Coordinators must also balance group dynamics, make exceptions for important individuals, and continuously replan when last-minute changes arise. Coordination is thus both tedious and delicate, requiring continuous monitoring and re-planning while respecting social norms and relationships.
Generative AI now possesses several key capabilities that could help with coordination, including deep reasoning to evaluate constraints and produce plans (Guo et al., 2025; Yao et al., 2023b, a; Stechly et al., 2025), and API tool-calling to take concrete actions (Trivedi et al., 2024; Yao et al., 2024; Masterman et al., 2024). However, people do not yet trust AI to act proactively on their behalf for complex and delicate tasks like human coordination problems. AI systems are prone to hallucinations, often lack sufficient context to understand the problem (Zhang et al., 2025; Athaluri et al., 2023), and may mishandle delicate social interactions (Zhou et al., 2023; Wang et al., 2024). Most importantly, they lack an awareness of when to stop and escalate to human oversight. As a result, current AI interfaces largely wait for users to issue prompts and micromanage responses. This limits productivity and prevents AI from relieving people of burdensome procedural work. To unlock proactive AI, we need frameworks that let humans observe, guide, and calibrate trust in AI over time.
Trust is not established by statements alone—it develops through repeated reliable behavior over time (Razin and Feigh, 2024; Fouragnan et al., 2013). Game theory suggests that trust between humans is built by repeated observation of behavior in varied circumstances (Kollock, 1998). Additionally, in the AI trust and safety literature, Constitutional AI (Bai et al., 2022) is a paradigm that uses human-articulated policies (called a constitution) that can align an AI’s behavior with human expectations and values. Policies in making expectations explicit - so that humans and AI have a shared understanding of the expected behavior. Repeated interactions and clear policies are two pillars of establishing trust that we build upon to help build trust in proactive AI.
Figure 1. System diagram of DoubleAgents illustrating a day-by-day ReAct workflow that couples policy-guided planning with human oversight and LLM simulation. Inputs (top left) include the organizer’s goal, seminar slots, speaker personas, and a toolbox of callable functions. The coordination agent (center) iteratively: summarizes state; selects applicable policies; proposes a plan and action that the user can Regenerate or Approve; executes by drafting and sending emails, or waiting for responses. If a reply falls outside policy coverage, issueFlag escalates an edge case to the user for clarification before proceeding. The context management (right) maintains time information, past plans/actions, email logs, user preferences, and persona data, continuously updating to ground subsequent steps. Simulated respondent agents (bottom right) generate realistic, persona-consistent behaviors and replies that drive the loop across days. For a detailed walkthrough, please refer to Section 4.
[…]
Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
We propose to build trust in proactive AI through simulation and policy refinement. Human behavior can be unpredictable, but AI-driven simulations can generate a wide range of realistic scenarios (Park et al., 2023), including late, incomplete, or uncertain replies. These allow coordinators to test and refine policies in safe environments. Policies specify how the AI should act (e.g., when to follow up, when to stop, how to phrase outreach) and when to escalate to human oversight. Hence, we propose to use simulation to expose AI behavior across repeated interactions, letting users correct and evolve policies iteratively. Visualizations of system state—goal progress, a timeline view, and person-level interaction histories—provide transparency and help humans intervene efficiently when needed.
We present DoubleAgents, shown in Figure 1, an interactive agentic system designed to automate aspects of human coordination while supporting user oversight and trust-building. DoubleAgents integrates generative AI agents in two complementary roles: (1) a ReAct Loop (Yao et al., 2023b) with AI planning and execution agents that can call tools to get appropriate context, query the current state of the system, and execute actions such as emailing potential speakers. This loop is extended to provide interaction for human interactivity and oversight at every step. (2) A LLM-based simulation of respondents that generate the varied and realistic human responses for various respondent personas (e.g. a professor who is very slow to respond, or a speaker who won’t commit). The LLM simulations allow users to experience repeated execution of the system to understand how it works, evolve the policies over time to suit their preferences, and potentially build trust that the system will act according to their wishes. By combining policy generation and refinement, action execution, and multi-agent LLM simulation within a user-centered feedback process, DoubleAgents helps users delegate repetitive aspects of coordination while retaining control of high-stakes interactions.
In a lab study (n=10) of people using the system with simulated responses over two study sessions, participants began hesitant to delegate but, through visible state, explicit reasoning based from policies, strong intervention points, and rehearsal on realistic cases, grew more comfortable relying on the agent for both planning and execution. In two real-world deployments of the system, we found that organizers adopted DoubleAgents as transparency, policy controls, and edge‑case flagging built trust; however, required more upfront context setup and tighter email integration.
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic workflow. Specifically, we contribute:
  • A policy-driven, intervention‑friendly approach to govern proactive AI behavior in an agentic and integactive ReAct framework to build trust with users.
  • An AI-driven simulation framework of human responses to support repeated observations of reliable behavior over time to test and tune policies for AI to follow.
  • A user study (n=10) and a deployment with two users to evaluate how users perceive and gradually trust proactive AI for a human coordination task of scheduling a speaker seminar.
2. Related Works
2.1. Agentic AI in Task Planning and Execution
“AI Agents” typically refers to AI that can perform tasks for people. This can be done through tool use (namely, calling an API) (Schick et al., 2023), or operating a computer (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). Although AI Agents show impressive potential to act, trusting agents to act on behalf of a human is a critical missing piece (Afroogh et al., 2024). The delegation of control to AI raises questions not only about competence, but also about alignment, transparency, and user oversight.
Planning and execution remain central yet unresolved challenges in the development of AI agents. A growing body of work highlights both the difficulty and the importance of task planning, as seen in benchmarks and simulation environments such as Tau-Bench (Yao et al., 2024), NaturalPlan (Zheng et al., 2024), TravelPlanner (Xie et al., 2024), and AppWorld (Trivedi et al., 2024). These studies underscore that robust, goal-directed planning, particularly in open-ended or multi-step domains, is not only technically demanding but also essential for delivering meaningful user value.
To address this challenge, recent work has explored how LLMs can scaffold user planning by decomposing high-level goals into actionable steps (Zhang et al., 2024). Techniques such as ReAct (Yao et al., 2023b), which interleave reasoning and acting, have gained traction in multi-step environments for improving agent adaptability and responsiveness. These approaches are often embedded within multi-agent systems to support domains like travel, scheduling, and personal task management (Lal et al., 2023). However, autonomous systems such as AutoGPT (Significant Gravitas, [n. d.]) often fall short in transparency, adaptability, and robustness, limiting their usability in real-world settings (Zheng et al., 2024; Yao et al., 2024). Furthermore, prior cognitive science research shows that humans rarely engage in exhaustive upfront planning due to cognitive load (Krieger et al., 2009; Teevan et al., 2016), highlighting the need for AI systems that can support incremental, iterative, and contextualized planning over time. Additionally, recent studies suggest that human–AI co-planning, particularly when agents expose their intermediate reasoning and allow users to guide or intervene, can improve both task performance and user trust (He et al., 2025). When users can monitor and influence agent decisions at each step, they are better able to understand, calibrate, and ultimately trust the system, especially in domains requiring high-stakes delegation or multi-party coordination.
2.2. Human Simulations with LLMs
Recent work has explored the use of LLMs in multi-agent simulations  (Park et al., 2023), revealing their surprising capacity to model (Horton, 2023), predict, and even exhibit human-like behavior - including taking actions they were not explicitly instructed to perform in order to achieve high-level goals such as forming routines, organizing events, and collaborating spontaneously (Park et al., 2022, 2023). These behaviors often mirror human reasoning and coordination strategies. Other studies have shown that these behaviors closely align with empirical observations of humans across various domains, including economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023; Horton, 2023), U.S. Supreme Court judgments (Hamilton, 2023), and over 80% of published and unpublished psychology experiments (Hewitt et al., 2024). These results suggest that LLM agents can serve as realistic proxies for humans in testing coordination, negotiation, and reasoning behaviors at scale.
AI systems that silate human behavior have already shown a striking capability to engage in complex, social behaviors competitively well with humans, for real world purposes. . Beyond simulation, AI systems like CICERO (Team et al., 2022) have shown how language agents can engage directly with humans in dynamic, multi-party settings. In online Diplomacy games, CICERO achieved a top 10% ranking among human players - illustrating the potential for LLM-based systems to engage in trust-sensitive planning, social reasoning, and strategic communication. The MATRIX framework (Pang et al., 2024) improved alignment in constitutional AI evaluations by useing simulated AI dialogues to anticipate a range of perspectives and surface important value trade-offs. Many systems have successfully used LLM red teaming  (Purpura et al., 2025) to identify potential threats in systems. Although AI simulations might not perfectly replicate individual user behaviors (yet) - they can be used to explore edge cases, probe systems, and engage in sensitive dialogs with humans.
2.3. Constitutional AI
Constitutional AI is a framework for aligning AI systems with human values by training models to self-critique and revise their outputs based on a set of guiding principles, or a ”constitution”(Bai et al., 2022). Unlike reinforcement learning with human feedback (RLHF)(Ouyang et al., 2022), which requires human evaluation for each model response, Constitutional AI allows the model to evaluate its own behavior using pre-defined rules. These constitutions are typically expressed as human-readable policies such as “Don’t promote illegal activity” or “Be helpful and respectful,” and can be authored by anyone. Building on this idea, Collective Constitutional AI (Huang et al., 2024) introduced a participatory approach in which over 1,000 people contributed rules to a shared constitution. Models trained on this crowd-sourced constitution demonstrated qualitatively different behaviors compared to those trained on developer-authored rules—highlighting the diversity of human values and community norms, and the implications this has for model alignment. To create constitutions, interactive systems like ConstitutionMaker (Petridis et al., 2024) have shown that LLMs can help transform users’ natural language feedback into generalized alignment policies. By providing real-time AI assistance in policy creation, such systems help users articulate their values more effectively, reduce cognitive burden, and make the process of shaping AI behavior more accessible and inclusive. These advances open up opportunities for systems that incorporate user-driven policy creation and alignment into workflows that have real-world impacts.
3. Seminar Planning: A Human Coordination Problem
Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate. Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time. Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—can erode trust and derail collective progress. These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
We study these challenges through one concrete and consequential setting: organizing an academic seminar series. A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term. Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own, such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
However, the most unpredictable and challenging part is dealing with responses. Some people do not reply at all and require multiple follow-ups to elicit a response. Some give only vague and tentative replies that need repeated confirmation. Some can be flakey and unreliable, offering slots but later canceling. And many unexpected problem cases arise: people might be on parental leave, ask for visa letters or funding support, recommend a different speaker, request a different date or talk modality, or raise logistical demands. In such human coordination tasks, you can rarely predict what is coming. Therefore, conventional scheduling tools such as Calendly [^1] https://calendly.com/ or When2Meet [^2] https://www.when2meet.com/ are ill-suited to this setting. They assume stable constraints and one-shot coordination, yet seminar scheduling requires ongoing negotiation, persuasion, and relationship maintenance. Unlike ordinary meeting scheduling, seminar speakers are doing you a favor, which means you cannot simply send them a link to pick a time—you must persuade and negotiate. Social dynamics and etiquette also play a large role: organizers must time their emails carefully to avoid overwhelming, annoying, or losing momentum with potential speakers, who are often busy scholars. Knowing when to nudge, follow up, or wait is subtle. Email remains the standard medium for professional and academic outreach, allowing organizers to adapt tone—formal, casual, or enthusiastic—to different speakers in ways scheduling links cannot. It also supports sharing rich context, such as talk topics, background, past speakers, or travel plans—far beyond simply choosing a calendar slot.
Thus, experienced organizers adopt personal strategies to cope with such a demanding task. For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again. Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints. When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption. For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability. Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate. An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
4. DoubleAgents System
Figure 2. A screenshot of DoubleAgents coordinating the assignment of four speakers to four seminar slots. (A) Policy Panel – Displays coordination policies that reflect user values and preferences. Users can add, edit, or delete policies, which guide the reasoning and planning of the coordination agent. (B) Interactive Chat Interface – Shows the full history of plans and actions, allowing users to monitor planning and execution, confirm or modify agent suggestions, and intervene when necessary. (C) Plan and Action Pop-up – A focused view within the chat history where users can review, edit, regenerate, or approve individual plans and actions. (D) Assignment Tracker – Visualizes tentative speaker-to-slot assignments and indicates each speaker’s availability. Dark purple denotes confirmed availability; light purple signals ambiguous availability requiring user confirmation. (E) Communication History Viewer – Displays the messaging history between the coordination agent and each speaker, organized by date and interaction. (F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
We present DoubleAgents, an interactive agentic system designed to investigate how users build and calibrate trust in proactive AI systems. Built on the ReAct framework (Yao et al., 2023b), DoubleAgents enables iterative planning and execution through a team of AI agents that assist users in complex coordination tasks. In this work, we focus on the use case of assigning seminar speakers to available time slots. The system takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots.”) and produces as output a finalized assignment of speakers to slots. DoubleAgents consists of six core components, grouped into the following three functional categories. First, it contains the Core ReAct Agents that carry out the reasoning, planning, and simulation logic that drive the coordination process:
  • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
  • Simulated Respondents - Represent background agents that emulate human respondents’ behavior by responding to user-generated emails based on predefined personas and descriptions.
Second, DoubleAgents incorporates the System Orchestration Modules, which manage the internal context and ensure robustness during decision-making:
  • Context Management Module - Maintains all relevant contextual information related to the task, including simulated respondents’ details (e.g., personas and descriptions), email logs from prior time steps, and the history of user-approved plans, actions, and emails.
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions.
Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can confirm or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
  • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.
In this section, we provide a walkthrough example to illustrate how DoubleAgents operates in practice. We also describe the implementation details of DoubleAgents. Consider the example of Lee, a PhD student organizing the HCI seminar series at his university. Lee plans to invite four speakers–Prof. A (a famous professor), Prof. B (a rising star professor), Student X (a senior PhD student), and Student Y (a PhD student who is a friend of Lee) – to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
Lee begins at the landing page, where he specifies his goals to the system. He types: “I am trying to organize a seminar with four speakers and four slots, help me reach out”. He then provides additional details, including the seminar location, and the dates of the four time slots. Next, Lee enters persona information for each speaker he wants to invite, which the system uses to generate simulated speaker profiles. The landing page is shown in Figure 3. To further customize communication, Lee uploads five of his past emails, allowing the system to adapt its generated messages to his style; for instance, casual for friends and formal for faculty. Finally, Lee initiates the simulation by clicking the “Start Planning” button, prompting the system to set up the scenario with the specified speakers and time slots. At this point, he is also shown a list of preset policies (see Figure 2(A)), which serve as guidelines for the coordination agent to begin planning. Lee reviews the policies, all of which appear reasonable. He may modify existing policies or introduce new ones to better guide the agent’s planning process.
Figure 3. The landing page to specify the user goals, seminar details, and the persona info for the speakers.
4.1. Context Management Module
At the beginning of the scheduling process, all information provided by Lee on the landing page is compiled into a structured problem context that the system consults at each step to solve the assignment task. DoubleAgents represents this problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history, etc.), and each value contains the corresponding content. As shown in Figure 1, an example snapshot of the problem context includes keys such as ”time info”, which reflects the seminar slots specified by Lee. The context is dynamically updated throughout the interaction, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
4.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 2(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 2(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix A.4.
4.3. Coordination Agent
4.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix A.1.1.
4.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, five experts were recruited to brainstorm and define a framework of rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 2(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix A.1.2.
(a) The pop-up window for a generated plan.
(b) The pop-up window for generated actions following the plan.
(c) The pop-up window for a generated email.
Figure 4. Examples for plan and action generation.
4.3.3. Plan Generation
Guided by this policy, the coordination agent generates the following plan (shown in Figure 4(a)) for Lee: “Send email to Prof. A, Prof. B, Student X, and Student Y to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”. Lee considers the plan and the rationale logical, since scheduling cannot proceed without first collecting availability information, and clicks ’Approve & Continue’ to move forward. The plan and its rationale are automatically recorded in the interactive chat interface (shown in Figure 1(B)), allowing Lee to review past activities and understand the rationale behind the plan. If Lee is dissatisfied with the proposed plan, he can either request the system to regenerate an alternative or directly edit the current plan himself, as shown in Figure 4(a).
4.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 4(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix A.1.3.
4.3.5. Action Execution
Following approval, the coordination agent proceeds to instantiate each action by generating email drafts for the designated speakers, including both subject lines and message bodies (Figure 4(c)). This is done by calling the email writing tool, as specified in the tool box shown in Figure 1. Specifically, it calls the writeEmail() API, which drafts emails given the current problem state, seminar details, user-provided templates, and historical email edits. These elements allow the system to learn and adapt to the Lee’s preferred communication style, capturing tone, level of formality, and domain-specific phrasing, when generating new emails. Note that email drafting is performed using GPT‑4o, which is invoked by the coordination agent during the execution of the writeEmail() tool.
Lee reviews each draft and identifies one instance where the tone is not appropriate—the email to Prof. A, a distinguished professor, feels overly informal. To address this, Lee types: “Please make the email sound more formal, as Prof. A is a distinguished professor I’d like to invite.” The system then regenerates a more suitable version. Lee further refines the message by editing the first paragraph to better reflect his personal voice: “My name is Lee, a PhD student in Computer Science. Your work in [XXX] has greatly inspired us, and I would be honored to invite you to speak at our HCI seminar series.” After reviewing and finalizing all drafts, Lee is satisfied and clicks ‘Approve & Send’, prompting the system to dispatch the emails to all invited speakers. This is accomplished by invoking the sendEmail() API by the coordination agent, targeting the appropriate speaker email addresses specified in the problem context. In addition, all email-related activities are logged in the dashboard (Figure 1(E)), enabling Lee to easily monitor the status and history of interactions with each speaker. Besides writing and sending emails, the coordination agent may invoke actions involving delayed responses, such as waiting for a speaker’s reply. This is achieved by calling the waitResponse() tool. In such cases, no message is generated or sent; instead, the agent pauses execution and either proceeds to the next pending action or advances to the next simulated day. Overall, we use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
4.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix A.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
4.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 1), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix A.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 5(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
(a) Email sent by the simulated Prof. A.
(b) Edge case detection given response from Prof. B asking about Zoom options.
Figure 5. The example email sent by the simulated respondents and the system message when detecting edge case.
4.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix A.2).
4.6. Dashboard Visualization
To support user understanding and foster trust, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 2(D), (E), and (F), which we describe in detail below:
  • Assignment Grid and Progress Tracker (Figure 2D): This panel visualizes tentative and confirmed speaker-to-slot assignments. It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled. This layer helps Lee assess how close the current plan is to completion at a glance.
  • Calendar and Timelines (Figure 2F): Anchored in simulated days, this temporal layer helps Lee reason about timing, deadlines, and follow-up actions. The calendar highlights both the current simulated day and the seminar dates, allowing Lee to prioritize actions based on urgency and scheduling constraints.
  • Communication Visualization (Figure 2E): This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status. It enables Lee to track ongoing communications, detect gaps, and assess follow-up needs.
These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time. Throughout the process, Lee interacts with the coordination agent and the respondent agents to collaboratively establish speaker-to-slot assignments for the seminar. Together, these visualization features provide Lee with a clear overview of assignment progress and a sense of what needs to be done on the current and subsequent days.
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix A.5. As shown in Figure 2(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
5. Technical Evaluation
DoubleAgents enables users to engage in iterative task planning and action execution through two core functionalities: (1) policy selection, which guides the generation of plans aligned with predefined coordination strategies; and (2) robust edge case detection, which identifies problematic or ambiguous situations and escalates them to the user for clarification and resolution. In this section, we describe the technical evaluation of the two components.
5.1. Policy Selection Experiments
5.1.1. Experiment settings
Policy selection refers to the process of identifying relevant coordination policies that guide the generation of high-quality, context-appropriate plans. Specifically, given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module is expected to retrieve all applicable policies that can serve as guidelines for the next planning step. For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts. To inform policy selection, we use the output of the getProgressSummary() API, which provides a high-level snapshot of the current task state. An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07. All other speakers are still pending. Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix B. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.
5.1.2. Results and Findings
Prompting
Techniques
F1 Score
Mean SD
Speaker-only .27 .054
Slot-only .11 .038
Slot + Speaker .27 .017
Summary-based .70*** .035
Table 1. The technical evaluation results for Policy Section comparing four prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state. The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection. Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval. Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
5.2. Edge Case Detection Experiments
5.2.1. Experiment settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix B.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix B.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
5.2.2. Results and Findings
Prompting Techniques
Accuracy
Mean SD
Zero-shot .79 .021
Few-shot w/o Policy .88† .000
Few-shot w Policy .93** .015
Table 2. The technical evaluation results for Edge Case Detection comparing different prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
6. User Study
6.1. Goal and Setup
To investigate the user-perceived effectiveness of DoubleAgents’ features and the dynamics of trust in human-AI collaboration, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their reliance on the systems, and their trust and comfort in using AI to automate tasks had shifted. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to their preferences. Participants completed a questionnaire assessing their trust in 15 different system features at three points: before the first session, after the first session, and after the second session. This allowed us to examine whether and when their trust and perceptions of the system changed over time.
6.2. Participants
We recruited 10 participants (7 female, 3 male; aged 21–30, M = 25.7) from our university community. All participants (10 of 10) reported using generative AI automation tools daily. Participants primarily used tools such as AI writing assistants (10 of 10), AI coding assistants (7 of 10), and AI design/creative tools (5 of 10). All participants reported solid understanding of generative AI: 7 described having in-depth knowledge of models and tools, while 3 reported a good general understanding of key concepts. Participants’ experience with human coordination tasks varied widely. 2 participants had moderate to extensive experience, having led team schedules or group planning multiple times. 3 participants had some experience, having organized small tasks or events independently. 4 participants had minimal experience, primarily helping or shadowing others in coordination tasks. Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
6.3. Procedures and Data Analysis
Test Scenario Preparation.
We select eight distinct personas identified by three organizing experts, representing realistic and diverse roles, research areas, and backgrounds (e.g., a fifth-year PhD student studying VR currently on the job market; a busy junior professor in NLP who spends much of her time on her startup). For each persona, we specify reply behaviors (responsiveness, email writing style; e.g., slow responder, writes only short emails) and optional issues (potential concerns they might raise with speakers; e.g., requesting additional funding even though their in-person visit is already funded). We split them into two groups and created seminar organizing instances. We set the timeline as a monthly seminar schedule, with the first seminar one month after the simulation start date.
Day 1, Preparation and Setup (30min)
Participants first provided informed consent and completed a short AI perception survey assessing their perceptions and experiences with AI automation tools, human coordination tasks, and demographics. They also edited a few seeded templated seminars organizing emails according to their tone to feed the system their writing style. They were then introduced to the DoubleAgents interface and task. One researcher gave a brief walkthrough using a visual diagram and demonstration, covering the system’s core components, including the planning-execution orchestrator agent loo, respondent agent simulation, policy and context modules, visualization, and the flag mechanism.
Day 1, First System Usage and Evaluation (60min)
Participants then completed a seminar scheduling task (40 minutes), interacting with four simulated speakers with diverse personas, response behaviors, constraints, and coordination tensions. During the task, participants articulated and shared their mental models and usage strategies for DoubleAgents. Afterward, they completed another time of AI perception survey and a system experience survey assessing experience, workload, and attitudes toward specific components such as control, usefulness, and trust. Survey questions were adapted from NASA TLX (Hart, 2006) and prior AI perception and long-term evaluation studies (Ross et al., 2023; Long et al., 2024). Finally, a short interview asked participants to reflect on their overall experience, including any insights or policies they would like the system to incorporate before Day 2.
Day 2, Second System Usage and Evaluation (60min)
Before users’ next DoubleAgent usages, the researcher first ensured that the system incorporated participants’ first-day planning context and any preferences they had established via policies or system edits. Participants then completed another seminar scheduling task (40 minutes), followed by the same AI perception survey, the system experience survey, and a short interview. Day 2 interviews included follow-up questions comparing participants’ mental models, perceptions, experiences, and expectations of the system between Day 1 and Day 2.
Data Collection and Analysis
We logged participants’ interactions with AI suggestions, including regenerations, rejections, and edits, which were incorporated into DoubleAgents’ implicit learning module and our analysis. Interview data were analyzed using thematic analysis to understand how users perceive the usefulness and trust drive about each feature, and quantitative data were analyzed using paired-samples t-tests to examine whether participants’ experiences and perceptions of the DoubleAgents and AI tools across usage and days.
6.4. Result: Users’ Perceptions in AI Systems Shifted After Usages
Figure 6. Average user perception scores toward proactive AI and comfort with automating the planning stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$; * at $p<0.05$; ** at $p<0.01$; and *** at $p<0.005$.
Increasing Trust and Desire for Proactive Control.
Figure 6(a) shows statistically significant changes in participants’ trust in AI systems to act on their behalf, from 3.20/7 (before using the system, based on their previous AI experiences), to 5.00/7 after two times of using DoubleAgents. P1 noted that even though DoubleAgents represented just one specific experience with a particular AI system, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.
Additionally, there was a significant increase in wanting detailed explanations before the system acts, from 5.00/7 (before) to 6.00/7 (after). This suggests that seeing explanations from DoubleAgents made similar explanations from AI systems more desirable. P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency, [which] I think is pretty good… if I didn’t get that, I would feel lost [and] not know what’s going on…
Figure 7. Average user perception scores toward comfort with automating the execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$ and * at $p<0.05$.
Increasing Comfort with Automating the Planning Stage.
Figure 6(b) shows participants’ attitudes towards proactive AI in planning became significantly more positive. All of these metrics showed statistically significant differences: they felt more comfortable letting AI visualize the current status (4.80/7 to 6.50/7), figure out potential time assignments in the planning (5.10/7 to 6.38/7), decide daily next steps (4.70/7 to 6.50/7), and act as an assistant to monitor or flag issues (5.50/7 to 6.50/7). For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker. She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.
For letting AI help monitor and point out potential issues, P1 described flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…
Increasing Comfort with Automating the Execution Stage.
Figure 7 shows that participants’ attitudes toward proactive AI in execution also became significantly more positive. For the tasks involving only drafting emails, as noted in the left column, all attitudes significantly improved, with comfort levels starting high and rising across the four types of emails (from around 6.0 before to nearly 7.0/7 after): initial outreach, follow-ups, clarification, and final confirmation. For the right column, draft-and-send tasks on users’ behalf, comfort also increased over time—rising for initial outreach from around 3.0/7 before to about 4.5/7 after, as well as for follow-ups and final confirmation. Participants demonstrated this after seeing DoubleAgents’ consistent performance over sessions: “I think [the system] acted quite reliably across two days, similar in terms of how it drafts emails [to the speakers], how it organizes language…” (P5). Participants also appreciated that the emails became more aligned to their tone and style over time—their email edits carried over to subsequent messages. For example, P4 mentioned that the system was learning her email writing style and that it remembered the phrasing she favored in email starters: “Compared to last time, I edited fewer times this time.,” which contributed to their trust in offloading more execution tasks to the AI, as well as a significantly increased desire for a system that aligns with and adapts to their styles over time (6.10/7 to 7.00/7, in Figure 6a).
One notable exception is drafting-and-sending clarification emails without approval, which showed little change (3.00/7 to 3.90/7 to 3.90/7). P10 explained that clarification emails often require nuanced, user-specific context or subjective judgment: “For clarification, I might want to double-check… because clarification sort of means that AI is also encountering a situation that was usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI… AI is already very good at pretending like it knows what it is talking about…” As a result, they preferred to remain in the loop, reviewing or guiding these clarification messages before sending.
Additional Trust Driver: Simulation
All 10 participants emphasized that the run-throughs of simulated test cases helped establish confidence in the AI system’s robustness, contributing to trust. They found the simulations useful for surfacing a variety of problems that were realistic and representative of real-life cases. For example, P3 mentioned “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule — it forced me to think about how I would handle that [in real life], and that was good. I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
7. Deployment Study
As the in-lab study demonstrated, the simulation was helpful, and participants felt comfortable trusting AI in planning and executing human coordination tasks. However, would real-world seminar organizers use DoubleAgents? If so, how would they use it—would they plan, draft, and send invitation, follow-up, or clarification emails in different ways? Would trust be established in a similar manner? What additional challenges might arise in real-world use?
To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited two organizers who were managing actual events during the preparation of this work. The goal was to observe how the system functions in realistic settings—including scheduling, tracking, email communication, and adaptive learning—and to understand how comfortable organizers are when the system sends real emails on their behalf.
7.1. Deployment Users and Procedures
Participants
We recruited two experienced seminar organizers (1 male, 1 female; average age 31), both with a good general understanding of generative AI and experience using AI for work—one on a daily basis. They are PhD students who have experience organizing research seminars for their departments. In this deployment, each is responsible for managing speaker events with audiences of over 30 participants: D1 is coordinating the Fall 2025 HCI seminar series at a university, selecting 10 speakers from a potential pool of 23 to fill 10 slots. D2 is organizing a panel discussion for a conference workshop, inviting 5–6 panelists from a potential 18 to fill a single slot. Each organizer was introduced to the system and used it to plan and manage their events.
Introduction, Setup, and Simulation Rounds
Organizers were first informed about the study and provided consent. They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops and trust mechanisms, illustrated with a simulation example from the user study. Organizers also completed the initial survey covering AI perception and demographics.
To set up the system, organizers prepared a realistic start sheet by replacing simulated data with real organizing data (real potential speakers’ names, first with fake email addresses and later with the real ones) and provided their Gmail SMTP credentials. Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw. Before connecting their real email accounts, participants completed a 2-hour simulation round using pre-generated speaker personas based on the real potential speakers, allowing them to safely familiarize themselves with the system’s workflows. Finally, they tested the real email-sending module twice to confirm that emails were delivered correctly.
Deployment
After the simulation familiarization, organizers used DoubleAgents to initiate the coordination process and track it for over a few days. They continued receiving system updates over the following week. We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning. Organizers’ system interactions—including email sends, scheduling updates, and system adaptations—were tracked. Qualitative feedback was collected via check-in meetings, focusing on usability, trust, concerns, and code support, while quantitative system logs were collected for system learning over the week.
7.2. Deployment Results
DoubleAgents’ Perceived Usefulness and Relevance in Live Deployment.
After the demonstration and simulation run-through, both organizers (D1 and D2) reported that DoubleAgents was highly relevant to their seminar-organizing workflow. They highlighted that the system fit seamlessly into their email-heavy tasks and that control settings and edge-case flagging enabled handling complex coordination scenarios that conventional scheduling tools like Calendly cannot manage. D1 emphasized that crafting emails is a delicate and important task, as it involves personalizing messages in order to “make the speakers feel they are special,” even when requesting a favor. Both D1 and D2 found DoubleAgents capable of generating good starter emails for edits at the beginning, as well as producing an acceptable final email to send. Organizers specifically appreciated that the simulation included the names of the real speakers and incorporated aspects of their personalities in constructing the respondent simulations and edge cases, which helped them anticipate challenges in live deployment.
System Onboarding, Testing, and Understanding in Live Deployment.
D1’s deployment setup took four days and five meetings. The first two meetings were devoted to debugging and ensuring the system ran smoothly for a scenario involving more than 20 speakers, 10 slots, and substantial context information for each speaker. D1 modified their invitation and follow-up email templates more than four times. Additionally, D1 updated the seminar speakers’ descriptions to enable the email generation module to generate precise, high-quality emails that incorporated extensive context information. After hardening the system and refining our onboarding protocols, D2’s setup took two days and two meetings. D2’s application in panelists invitation were also simpler (e.g., shorter panelist descriptions, invitations for one-day panels), which contributed to the faster setup. Both D1 and D2 reported that the practice simulation rounds were valuable, and that sending testing emails through SMTP allowed them to gradually grasp the system’s capabilities and limitations. This illustrates that system setup and onboarding are consistent during deployment, while the required effort scales with task complexity and the amount of context data.
Specifically, organizers emphasized the importance of supplying sufficient context to generate highly personalized emails. D1 mentioned, “[context management] is really tricky.” For example, one speaker D1 tried to invite had previously corresponded with a different organizer in a prior semester about her seminar availability on Fridays in a separate email thread. This semester, the seminar had shifted to Thursdays. To manage the context from past email threads, schedule updates, and templates, D1 opted to start a new thread with the speaker, as DoubleAgents at the time worked with new threads rather than replying within existing Gmail threads. In another case, D1 had already discussed seminar opportunities with a speaker on Discord but initially omitted this context during the early simulations. When composing the email, they noticed that the system had selected a generic template, so they updated the backend context with the appropriate tone and information to create a suitable follow-up. After practicing with real-speaker data in simulation, D1 finalized a preferred template for live sending. During the actual sending, they continued to refine messages—adding details and small fillers—and observed that these edits were retained and applied to subsequent emails.
Email Initiation, Response Handling, and Edge Cases in Live Deployment.
During live deployment, D1 sent emails to nine speakers. Over the earlier simulation rounds, she carefully perfected her initial templates and preferred writing styles via the implicit style transfer modules. Despite this, she still spent significant time refining the first few real emails. For the remaining speakers, she adopted a common real-world practice: copying previous templates into the text editor and adapting them for each recipient, which sped up the process. She mentioned that “the system’s email suggestions are helpful, though for similar invitation messages, I don’t always need a new draft each time.” D1 also used the system’s built-in AI chat feature seven times during initial email sending for phrasing recommendations. She noted that this support reduced mental demand and felt similar to receiving quick email feedback from a co-organizer, a practice she commonly uses in real-life organizing. D2, on the other hand, relied on the system’s automated drafting and style adaptation, demonstrating that trust and efficiency increased with experience. Interestingly, for some longer emails requiring scrolling to read, D2 skipped reviewing the full content and proceeded directly, trusting the system’s style transfer. At the end of the invitation email sending day, D2 sent emails to five potential panelists, relying directly on system-generated drafts based on his perfected templates and implicit style transfers from the previous simulation rounds.
After the initial invitation email were sent, D1 received responses as quickly as 4 minutes, 1 hour, and 2 hours. As replies came in, organizers encountered more complex edge cases than those seen during simulation: all three responding speakers had overlapping availability on the same dates, creating interlocked scheduling conflicts. DoubleAgents logged the replies in the backend and suggested that D1 wait for the remaining six speakers while coordinating with the student speaker and the faculty member closest to her school to adjust availability. D1 instead opted to wait for three days before responding (one day more than the policy instructed), as she was occupied with other commitments and felt comfortable replying later. In contrast, D2 received three negative responses after a full day and did not encounter similar complications, and the system suggested waiting for the remaining two people to reply.
Trust Establishment in Live Deployment
Over time, D1 emphasized that trust depended on the system’s closeness to ground truth—specifically, whether emails were actually sent and whether responses were logged correctly and efficiently. Accurate alignment between the system’s state visualization and the real state of email threads was critical to their trust. Also, D1 noted that, over time, understanding the system’s capabilities and limitations helped her feel more comfortable using the tool. D2 reported that they appreciated having policies and rationale always visible and editable, in contrast to AI tools that hide their reasoning. After sending real emails, D2 remarked, “I trust AI for specific use cases more,” explaining that they found the system to be dedicated to and capable of generating high-quality email.
At the same time, D2 compared DoubleAgents with other tools they were familiar with, abstracting their trust establishment process and the underlying barriers. He noted that the email interface differed from Gmail, introducing a small trust barrier. They compared Gmail’s richer text editor, easier insertion of tables and visuals, and familiar thread visualization—where they know what to expect when clicking—to our interface. They also expressed concern about formatting consistency for recipients (e.g., indentation, font sizes) in emails sent via DoubleAgents’ Gmail backend, noting that such issues can occur even within Gmail phone apps, which made them unsure how messages would appear to recipients. This concern was addressed through several rounds of email formatting and link embedding tests, during which they sent mails to their personal Gmails and checked them in parallel. Regarding integration, D2 felt that strong state visualization and tracking supported their trust through natural mapping. They believed their current practice of using spreadsheets to track speaker invitation progress could be integrated into the system’s visualization. For example, he usually rank the lists to monitor progress, and instead of serving solely as a display, the spreadsheet also functions as a goal-setting and action-taking tool. Thus, he expressed interest in making DoubleAgents’ visualization more interactive and editable, allowing them to directly initiate system actions. For actual adoption, he emphasized the need for scalability and further automation, requesting dynamic fields features similar to other email client tools like MailChimp. They also compared real email SMTP sending to Zoom microphone testing, emphasizing the value of having a trial, closed, and safe environment with strong validation—such as sending emails to themselves first.
Trust Establishment Mechanism in DoubleAgents How Humans Build Trust With Humans Nielsen Heuristics (Nielsen, 1992) Human-AI Design Challenges (Yang et al., 2020)
(1) Consistent and robust
Policy-based conditional reasoning grounded in human social cognition; reliable handling of varied test suites and edge cases
Trust grows when others are reliable, predictable, and follow shared norms or values. Consistency over time signals dependability. (Lansing et al., 2023; Rempel et al., 1985; Mayer et al., 1995; Beratan, 2007) - Consistency and standards
- Error prevention
- Help users recognize, diagnose, and recover from errors
- Unpredictability of AI behavior: users need repeated, consistent outputs to form expectations
- Data dependency: robustness across edge cases builds confidence
(2) Controllable and aligned with user values
User can confirm and intervene at every step; system adapts to users’ editing history and evolving policies
Trust strengthens when people respect autonomy, allow mutual control and negotiation, and align with shared goals or values (Lansing et al., 2023; Kramer, 1999). - User control and freedom
- Flexibility and efficiency of use
- Conceptual misalignment: user oversight helps align system behavior with goals
- Evolving mental models: adaptation supports shifting user practices
(3) Understandable and explainable
Real-time state tracking and visualization; clear reasoning and a chat window to ask questions
Humans trust others who are transparent, communicate intentions, and explain their actions clearly (Lansing et al., 2023; Kramer, 1999). - Visibility of system status
- Help and documentation
- Match between system and the real world
- Explainability gap: reasoning and state tracking reduce opacity and support accurate mental models
Table 3. How DoubleAgents’ trust-building dimensions map to human–human trust building literature (Rempel et al., 1985; Mayer et al., 1995; Kramer, 1999; Lansing et al., 2023; Beratan, 2007), Nielsen heuristics (Nielsen, 1992), and human-AI interaction design challenges (Yang et al., 2020).
8. Discussion: Trust Establishment with AI Agents Over Time, Consistency, Control, and Understanding
DoubleAgents illustrates how users establish trust in a novel AI system over time. Users initially expressed low trust because they were unfamiliar with the system, but through exploration, testing, learning, and habituation, they gradually began to rely on it. This aligns with prior longitudinal studies on AI tools (Long et al., 2024). Understanding how this trust emerged sheds light on the mechanisms by which human–AI relationships can evolve.
As shown in Table 3, DoubleAgents’ trust-establishment mechanisms center on three interdependent qualities: consistency, control, and understanding. These dimensions align with patterns of human–human trust development, are supported by classic usability heuristics (Nielsen, 1992), and address well-known challenges in human–AI interaction design (Yang et al., 2020).
First, one main design goal of DoubleAgents is to be consistent and robust. The system employs policy-driven, conditional reasoning to stabilize its outputs and ground its behavior in shared human values and social norms. This design enables reliable performance even across diverse testing suites and edge cases, allowing users to form stable expectations—echoing how consistency signals dependability in human relationships and mitigating issues of unpredictability and data dependency in AI systems. Second, DoubleAgents incorporates mechanisms to remain controllable and aligned with user values. The system enforces stepwise confirmations and allows user intervention throughout planning and execution, while adaptively learning from users’ past editing histories and evolving policy additions. These mechanisms promote a sense of autonomy and mutual negotiation, reflecting how humans build trust through control-sharing, and addressing conceptual misalignments and evolving mental models in human–AI collaboration. Third, DoubleAgents is designed to be understandable and explainable. It offers real-time state tracking, clear visualizations, explicit reasoning for its actions, and a chat interface for user queries. These mechanisms increase system transparency and communicability, similar to how open communication fosters trust between humans, and reduce the explainability gap in AI systems.
Collectively, these design mechanisms enabled users to gradually transition from skepticism to reliance. By operationalizing consistency, control, and understanding, DoubleAgents supported the iterative habituation process through which users internalized the system’s behavior and incorporated it into their workflows. This suggests that trust in proactive AI systems is not a fixed property but an emergent outcome—shaped by interaction designs that allow users to explore, test, and progressively align the system with their own goals and values.
Several participants compared building trust in DoubleAgents to their prior experiences with coding automation tools such as Cursor and CLine, reflecting the CS background of most participants. For example, P10 described first trying the tool on a safe project in a controlled environment (not pushing to a live website) to see how well it worked, illustrating the role of consistent system performance in building trust. They also highlighted that Cursor offered strong control and state-reverting capabilities, which allowed them to confirm each step and explore outcomes safely. Over time, they learned how to craft better prompts, demonstrating alignment with their goals and values. Finally, participants emphasized the importance of understandable reasoning: both Cursor and DoubleAgents provide stepwise explanations for each action, which helped them make sense of the system’s behavior and develop confidence in its outputs. As P4 reflected: “Most of the trust I had with Cursor was because it would ask at each step and then explain its reasoning for that step—and it was accurate. The same thing here [in DoubleAgents]: it’s accurate, and each step provides reasoning. It’s from the frequent confirmation and the reasoning always being correct that I built trust.”
These accounts illustrate that consistent performance, controllable interactions, and clear explanations collectively scaffolded the sophisticated trust formation process between AI and users, aligning users’ experiences with the mechanisms designed into DoubleAgents and with broader theories of habituation and technology acceptance.
9. Limitations and Future Works
First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios. While these simulations were reviewed by domain experts and deemed realistic, they may still lack the richness of real-world settings. Future work could make the ReAct-based coordination agent more user-aligned by incorporating users’ past email or organizing data and modeling their personal communication styles and values. Similarly, the respondent agent simulation could be made more realistic by drawing on a more diverse set of real human behavioral data. Second, our system used different models for different tasks—o3 for reasoning tasks and 4o-mini for other tasks. Although our technical evaluation showed they performed well, future models with stronger social and human cognition capabilities, as well as incorporating users’ data, may achieve better planning and execution. Third, although we deployed the system with two real organizers in real life, we also consulted others whose scenarios introduced additional complexities where the system may not perform perfectly. For example, one seminar organizer mentioned theirs involved bilingual invitations, and others described co-organizing dynamics, both of which introduce challenges not addressed in our current design. Finally, all our participants were experienced users of AI tools. Their familiarity may have positively biased their trust and comfort levels. Future work should include participants with a broader range of AI experience to better understand adoption barriers among less experienced users.
10. Conclusion
We present DoubleAgents, an agentic planning system that incorporates clear human intervention points, rich state visualizations, explicit policies, and uncertainty/edge‑case flagging to automate human coordination. In a two‑session lab study (n=10) and a week‑long deployment with two organizers, participants reported significantly greater comfort delegating planning and execution to DoubleAgents. Simulated respondents provided a safe proving ground to observe policy‑aligned planning and consistent system performance. From these studies, we distill three trust‑by‑design mechanisms—consistency, control, and understanding—operationalized via policy‑driven reasoning, stepwise approvals, implicit system adaptive learning, and a real-time system state. These patterns generalize beyond seminar scheduling to proactive agents design that must act while remaining auditable and aligned.
Appendix A DoubleAgents Prompts
A.1. Coordination Agent
A.1.1. Progress Summary
You are an expert seminar organizer assistant. Your task is to analyze the current progress and return a JSON object. Input information may include: - Speaker details (persona and description) - Possible dates - Scheduling rules - Email log - History of approved plans/actions - Time context - Any user suggestions You should mainly look through Email log and history of approved plans/actions to analyze the current progress. Do not assume any additional actions have been taken. Your JSON output should contain: - summary: A concise summary of the current state, including which days are scheduled, what actions have been taken, and the last date we reached out to each speaker (for tracking and follow-up). - un_or_filled_slots: Two lists of dates — one for filled slots and one for unfilled slots. - pending_speakers: A list of speaker names who have not yet responded or are pending (how many days since we last reached out to them). When mentioning speakers, use the format “English Name (ID)”. Avoid vague terms like *all speakers* or *pending speakers*. ---------- [Task Rules] [Possible Dates] [Time Context] [Speaker Information] Here are the history of Approved Plans and Actions: [Approved Plans and Actions] -------------------------------- Here is the history of past user edits and implicit learning (recent, action edits only): [Implicit Learning] -------------------------------- Past Email Log: [Email Log]
A.1.2. Policy Selection
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy Return a JSON list of the indices of the applicable policies (e.g., [1,2]) or [1]. If none apply, return an empty list []. [Main Reasoning Result] [Coordination Policies] Return ONLY the JSON list of indices. [Implicit Learning/User Edit History (recent, action edits only)]
A.1.3. Action Generation
You are an expert seminar organizer assistant. Your task: Given the organizing progress summary, the selected coordination policies, and the current time context, generate a concise plan for what must be accomplished today only. Output format: Return only one single string that contains: 1. The plan (concise, no bullet points, no unnecessary multiple sentences). 2. A short rationale starting exactly with "Why this plan? ...". The rationale should be about 40 words explaining why this plan was chosen according to the task progress summary context and how the policies/time context informed the decision. Example: "Send email to X to request availability for slots XXX. Why this plan? As we are just beginning the scheduling process and have not reached out to X yet, we followed policy #6 (highlighted on the left) to obtain X’s availability for the slots. This will help jumpstart the process."" Guidelines: - Use each speaker’s real English name. - Focus strictly on today (ignore tomorrow or later). - Valid actions (choose strategically): 1. "Send email to X to request availability for slots XXX" (initial outreach). 2. "Follow up with X" (if email sent but no reply). 3. "Wait for response from X" (if email sent and still within expected waiting window). 4. "Confirm and assign slot for X" (only if X has shared availability; avoid finalizing too early unless time is urgent). 5. "Send clarification email to X" (if clarification on dates/details is needed). - If multiple actions are needed, combine them into one clear sentence (e.g., "Send email to X, Y, and Z to request availability for slots A, B, and C."). - Do not include any other tasks (e.g., slides, addresses, future prep). - Remember what day today is and when the last emailing was done (relevant for some policies). - Consider time urgency: - If few days remain, prioritize immediate action. - If weekend, default to waiting for responses unless a policy specifies otherwise. -------------------------- Time Context: [Time Context] Task Progress Summary: [Progress Summary] -------------------------- Applied Coordination Policies: [Applied Policy] -------------------------- Implicit Learning/User Edit History (recent, action edits only): [Implicit Learning - Action Edits] --------------------------
A.1.4. Email Drafting
I am [Organizer Name], a seminar/workshop organizer. Please help me write an email to [Speaker Name] about [Email Goal]. Please make sure it conveys that. Try to use a casual tone. If I haven’t given you any details, keep it concise and avoid unnecessary formalities. Sometimes it is a full email to invite some, while sometimes it might be just a followup email or a clarification / confirmation email. So you should write different emails according to the situation and past email history. Please return a JSON object with the enhanced email containing: - "from": organzier (just organizer, not names), - "to": recipient (ensure it’s just speaker’s Full name, no other information, dont change) - "subject": subject - "body": enhanced email body that is professional and clear Make sure the email is natural language, professional, and incorporates all context provided. Consider the time urgency when writing the email. If only a few days remain until the seminar dates, convey urgency appropriately. For follow-up emails (like following up / reminder emails -- no need to mention the seminar information.) You need to explicitly mention the dates, don’t say "next seminar/next Thursday" or "Nov 1, 8, 10". You need to list the dates individually. Start the greeting with "Dear [Speaker name]". --------------------------- Here are some information of all the speakers we want to invite Context: [Progress Summary] Past Email History with that Speaker: [Past Email History with Speaker Name] --------------------------- Time Context: [Time Context] --------------------------- Seminar Information (can be opted out if this is not the first email) [Seminar Information] --------------------------- Users’ template examples for invitation, if user has provided some, please try to follow them: User’s history invitation email : [Invitation Email Template] User’s history followup email: [Invitation Followup Template] User’s history confirmation emails : [Invitation Confirmation Template] --------------------------- When necessary, you should adapt some style edit that I have already taken on previous emails. However, be smart on whether they are the correct people to do so. [Implicit Learning - Email Edits]
A.1.5. Implicit User Preference Learning
Plan and Action Edits
You need to describe the user’s edits by comparing the AI/system {item_type} with the user-edited {item_type}, making use of the history context. Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the {item_type} when [context about the stage or progress of the allocation].’ Example: The user removed Tracy and Jack from the AI-generated outreach plan but kept others, such as Tanya and Adam, who are both faculty members. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
Email Edits
You need to describe the user’s email edits related to format and style by comparing the AI/system item_type with the user-edited item_type, making use of the history context. You need to first analyze if the user edits relate to format and style. If not, you should return "No learning needed as the user edits are not related to format and style." Note that you do not need to analyze the user edits if it relates to content (e.g. specific dates) Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the item_type when [context about the stage or progress of the allocation].’ Example: The user changed "Best regards," from the AI-generated outreach plan to "Thanks," in the email when initiating outreach to Jack regarding available dates. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
A.2. Edge Case Detection
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG: SpeakerName seems unavailable due to maternity leave and suggesting a lab substitute. Should I follow up with the proposed colleague or reassign the slot to another invited speaker? Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG: SpeakerName’s traveling during scheduled week requests swap to May. Should I reschedule them to May or retain current slot and reassign to another speaker? Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG: SpeakerName’s uncertain due to family reason, with final notice only one day prior. Should I secure a backup speaker now or wait for their last-minute confirmation? Here are the policies: [Policies]
A.3. Simulated Respondents
You are simulating a seminar speaker. Your persona: [Speaker Persona]. Name: [Speaker Name]. The organizer’s name is: [Organizer Name]. The current simulation day is: [Day ID]. You have received the following email thread (latest at the bottom): [Email Log] Possible dates to choose from: [Speaker Availability]] As the speaker, you can: - Reply to the organizer’s email, providing your availability, asking questions, or giving other relevant information. - If you have questions for the organizer, you can draft them in your reply and wait for a response. - You may choose to wait for more information before replying, if appropriate. - You can reply to multiple emails in the thread if you wish. - You may also choose to wait for a specific number of days before replying (e.g., "wait 2 days and reply on day 5"). If you choose to wait, specify the simulation day you intend to reply. Decide what you want to do next: - If you want to reply, write a detailed, persona-driven reply (possibly including questions for the organizer). - If you want to ask a question and wait for a response, draft your question(s) and indicate you are waiting. Do not ask questions that are too specific (’like equipment setup’ etc, this is a scheduling task, they only care if you are available on the date.) - If you want to wait and not reply yet, explain why, and specify the simulation day you intend to reply (e.g., "wait_until_day": 5). Return your action as one of: ’reply’, ’ask_question’, ’wait’. If you choose to wait, also return a "wait_until_day" field (integer simulation day to reply). Then, provide your reply email (if any) as a JSON object with keys: from, to, subject, body, timestamp. Remember, your subject should be just a short first-point-of-view summary of your reply (less than 10 words, like "I can make it on Sep 5", "I can’t make it on Sep 5", "Sorry I need a few days to think." etc). Also, provide a short "RATIONALE" section (2-4 sentences) explaining your reasoning and thought process for this turn. Format your response as: ACTION: <reply|ask_question|wait> WAIT_UNTIL_DAY: <integer or ’none’> EMAIL: <JSON object or ’none’> RATIONALE: <your reasoning> EMAIL should be like: "from": "XX", "to": "XX", "subject": "XXX", "body": "XXXXX", "timestamp": "XX" dont output json ‘‘‘ etc
A.4. Free-form QA Chat
You are an expert assistant. The following is an email log from a seminar simulation system. Answer the user’s question based on the email log. If the answer is not directly available, say so, but try to be helpful and concise. [EMAIL LOG] [USER QUESTION]
A.5. Daily Summary
Today is simulation day [Day ID]. Here is the email log for today (may include previous days): [Email Log] Please provide a concise 40-word summary of the progress made today, including any key actions taken, responses received, and remaining issues. The summary should be suitable as a notification for users in a sidebar. Dont include today’s date in the summary -- just summarize the progress.
Appendix B Technical Evaluation Details
B.1. Policy Selection
Here are the prompts used in the technical evaluations for policy selection:
B.1.1. Experiment prompts
speaker-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
summary-based
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Main reasoning result: {progress_summary} Coordination Policies: {policy_list} Return ONLY the JSON list of indices:
B.2. Edge Case Detection
B.2.1. Test suite construction
We employ GPT‑5 to generate a set of 10 edge-case (flaggable) test cases and 10 non-flagged (clear) test cases, with all examples validated for quality by the lead authors. To enhance linguistic and contextual diversity, each test case is further expanded into five stylistic variations, reflecting different tones and levels of formality. These variations are conditioned on five distinct speaker personas: a senior renowned faculty member, a junior faculty member, a senior PhD student, a junior PhD student, and a close peer researcher. In total, 100 test cases are constructed and used as our test suite for edge case detection.
B.2.2. Prompts to generate linguistic variations and test data
You are a helpful assistant. You need to generate test cases for me to test my flag detection assistant. I am organizing a seminar series, inviting speakers. You will be acting as the simulated speakers, writing response emails to my invitation email. You will be given cases where the system should mark as clear and that are not edge cases. For each case, you should generate exactly five different variations of responses in the role of the following five speaker personas: 1) a famous distinguished professor who is always busy 2) a junior faculty who is trying to establish herself in the field 3) a senior phd student who is able to graduate and is looking for a job 4) a junior phd student who is just starting his phd journey 5) a close friend of mine who is also a phd student researcher. The following are the cases where the system should mark as clear (i.e. labeled as "0"): Please generate the test cases in the following JSON format: """ { "case_description": $\{"persona":"persona\_description","response":"response","flag":"0$"\} } The case_description is the title of each scenario shown above. Please generate exactly five variations of the response for each case
B.2.3. Experiment prompts
zero-shot
You are an expert seminar organizer assistant. Analyze this response from a speaker invitation. Response: response You need to flag edge-case issues that may require user confirmation or special and urgent attention If an edge-case issue is found, simply output "FLAG", otherwise output "Clear". Your Output:
few_shot without policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Response: {response} Your Output:
few_shot with policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Here are the policies: - "Don’t be shy to ask for more availability if they don’t give you many. They will probably reply more if you beg." - "Some people don’t read or respond to emails right away because they’re busy. You should wait for their response, but it’s okay to follow up every two workday days if you haven’t heard back.", - "If someone hasn’t responded to the email after three reach-outs, then don’t contact them again.", - "Ambiguity is okay. People can typically do the days they are vague about — you can prioritize asking for more availability and just confirm the ambiguity right before assigning them at the end.", - "Some speakers—especially faculty—are more desirable to schedule. We should prioritize their availability by assigning slots to faculty over students and asking students to be more flexible or to provide additional dates to accommodate them.", - "If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots." Response: response Your Output: