### v2

# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people can align systems that act on their behalf with their goals, values, and situational expectations. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their use of agentic behavior before live deployment. We evaluate DoubleAgents in a two-day lab study (n=10), three deployment studies, and a technical evaluation. Results show that participants initially hesitated to delegate but used simulation to probe system behavior and adjust policies, gradually increasing delegation as agent actions became better aligned with their intentions and context. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that simulation helps users effectively manage real-world tasks with higher complexity and uncertainty. We contribute interactive simulation as a practical pathway for users to iteratively align and calibrate agentic systems.
generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; arxivbooktitle: Proceedings of the 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. DoubleAgents is an agentic AI coordination tool that uses interactive simulation to let users experience and align the full ReAct cycle before live deployment. Starting from an initial setup (left)—including a simulation goal, simulated test-case personas, and an initial set of policies, and email templates—users can undergo multiple simulated ReAct cycles (middle) to rehearse planning, review and approve actions, execute plans, detect new edge cases, and iteratively refine policies via stop hooks, and email templates over repeated test cases and time. This process results in (right) the final speaker-slot allocation, as well as a user-aligned set of policies, emails, and stop hooks before the system sends real emails.
1. Introduction
Figure 2. System diagram of DoubleAgents illustrating a day-by-day ReAct workflow that couples policy-guided planning with human oversight and LLM simulation. Inputs (top left) include the organizer’s goal, seminar slots, speaker personas, and a toolbox of callable functions. The coordination agent (center) iteratively: summarizes state; selects applicable policies; proposes a plan and action that the user can Regenerate or Approve; executes by drafting and sending emails, or waiting for responses. If a reply falls outside policy coverage, issueFlag escalates an edge case to the user for clarification before proceeding. The context management (right) maintains time information, past plans/actions, email logs, user preferences, and persona data, continuously updating to ground subsequent steps. Simulated respondent agents (bottom right) generate realistic, persona-consistent behaviors and replies that drive the loop across days. For a detailed walkthrough, please refer to Section 4.
Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
Human–AI alignment is one of the defining challenges of the 21st century, especially for agentic AI: systems that go beyond answering questions to taking actions on users’ behalf (Leike et al., 2018; Soares and Fallenstein, 2014; Hendrycks et al., 2021). Following the Bidirectional Human–AI Alignment framework defined by Shen et al. (2025), we view alignment in agentic settings as an ongoing co-adaptation process between users and systems. On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect an individual user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024). On the human side, users must continually develop appropriate mental models, usage patterns, and AI literacy so they can effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023). As agentic AI becomes more common, we need practical ways to assess and improve whether these systems remain aligned as they act in the world.
Today, many popular agentic use cases, such as agentic coding or “vibe coding”, are relatively low-stakes and operate in closed sandbox environments (Horikawa et al., 2025). In these domains, alignment is often easier because systems can leverage large-scale, well-structured public data (e.g., GitHub), and success criteria are comparatively legible (e.g., whether code compiles, passes tests, or produces the desired output) (Chen et al., 2021; Jimenez et al., 2024). However, in more complex tasks that must adhere to human and domain knowledge as well as social cognition and norms, such as writing emails to real people to request favors, where etiquette is highly context-dependent, agents cannot reliably align on their own (Afroogh et al., 2024; Gabriel, 2020). These tasks often require nuanced, case-by-case judgment that is difficult to infer from generic training data alone and may demand substantial human oversight if deployed naively.
One representative scenario is human coordination, such as scheduling workers with the right skills and experience for shifts they are available for or inviting researchers to a speaker series. In this work, we focus on the speaker-series scheduling setting. Organizers must secure a speaker for each slot (e.g., 4-5pm on Fridays), coordinate availability across multiple candidates, and handle uneven responsiveness (e.g., some invitees reply late or not at all). Responses can also introduce socially sensitive and context-dependent requests: “Can you reimburse travel?”, “Can I do Zoom?”, “I can’t make it on Fridays, what about Thursdays 4-5pm?” “Sorry for the last-minute cancellation—I now have a conflict,” or “My schedule is uncertain; can you follow up next month?” Determining the right next step is rarely optimal in a mathematical sense; it depends on the organizer’s values and preferences, which we refer to as “policies” (e.g., “Offer Zoom only if in-person is not feasible and no one else can do that slot,” “Prioritize senior faculty over PhD students when conflicts happen,” or “Follow up after two business days unless urgent”). In practice, because these “actions” are carried out almost entirely through email communication, the organizer must decide which email to send next (and to whom), how to phrase each email with the right content and right tone, and when to send it.
While an LLM-based agent may generate plausible actions and write acceptable emails, it often lacks the organizer’s specific norms and cannot reliably anticipate how preferences should change across cases. Alignment therefore must happen on both sides: the system needs ways to acquire and apply three user-specific automation refinements: (1) coordination policies that guide planning and action choices, (2) email templates and writing style that match the user’s voice and social intent for each action, and (3) stop hooks that define when the agents need to pause and request for user confirmation; and the user needs support to form a clear mental model and specify these three preferences, especially early on, when they may not yet know what rules, tone, or escalation boundaries they want. The central challenge lies in how can we accelerate this bidirectional alignment so the agent can act more autonomously over time, without requiring constant babysitting.
Our core approach is interactive simulation to support iterative automation refinement, enabling alignment between users and agents. As modern AI can simulate diverse human behaviors (Park et al., 2023, 2022), we use simulation as a testbed to expose users to realistic coordination scenarios, including problematic or edge-case behaviors, and to help them define and refine their communication strategies. During simulations, the agent proposes plans and actions (e.g., inviting multiple candidates, deciding when to wait versus follow up, or holding a slot for a high-priority speaker), and the user evaluates whether these choices are socially and contextually appropriate. This mirrors ideas from software engineering such as unit and integration testing: a system earns confidence not by succeeding on a single happy path, but by handling a range of representative cases. In our context, however, completing a task is not enough: agentic AI systems must be user-specific and value-aligned, reflecting personal priorities, norms, and preferences.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. DoubleAgents integrates generative AI to simulate two complementary roles, as Figure 2 shows. First, it uses a ReAct loop (Yao et al., 2023) that pairs coordination agents for planning and execution. The agents can call tools to retrieve relevant context, query the current task state, and execute actions such as drafting and sending emails to potential speakers. We extend the loop with explicit interaction checkpoints so users can review, intervene, and approve decisions throughout planning and execution. Second, DoubleAgents includes an LLM-based simulator of respondents that generates diverse, realistic replies for different personas (e.g., a professor who responds slowly or a respondent who introduces funding or logistics constraints). This simulation enables repeated run-throughs of difficult cases, helping users understand how the agent behaves, stress-test policies against edge cases, and iteratively refine preferences and autonomy boundaries. Across interactive simulation runs, the primary outputs are not only a candidate schedule (speaker–slot allocation), but also a reusable set of user-specific alignment artifacts: refined policies, updated email templates, and calibrated stop hooks that can be applied when the system is deployed to send real emails.
We evaluate DoubleAgents through a technical evaluation, a two-day lab study (n=10) in which participants used DoubleAgents with simulated respondents, and three real-world deployments. In the lab study, participants initially hesitated to delegate. However, over time, through multiple simulation rounds and the two study sessions, participants increasingly delegated, as visible state, policy-grounded reasoning, and clear intervention points (including stop-hook flagging) helped them calibrate autonomy and build confidence in the agent’s planning and execution. By the end, participants also produced reusable policies, stop hooks, and email templates aligned with their voice and style. Three week-long deployment studies demonstrate DoubleAgents’ usefulness and relevance for real-world, complex human coordination tasks. DoubleAgents was connected to organizers’ real email workflows and sending emails to real invitees, as well as incorporated incoming real speakers’ replies into its ongoing coordination state. Organizers reported that interactive simulation beforehand helped them prepare for deployment by crafting strong initial policies and email templates for real-world use.
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic AI workflow. Specifically, we contribute:
  • A design perspective on alignment in agentic AI, realized through a policy-driven, intervention-friendly ReAct framework that supports transparency, human oversight, and alignment with human goals and values.
  • An interactive simulation mechanism that models realistic human responses, enabling users to safely explore agent behavior, understand system dynamics, rehearse interventions, and iteratively refine policies, email references, and stop hooks before real-world application.
  • Empirical results from a user study (n=10) and three real-world deployments, examining how interactive simulation and policy refinement support alignment in agentic AI for human coordination.
2. Related Works
2.1. Alignment in Human-AI Systems
As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge. A key strategy is reinforcement learning from human feedback (RLHF), which fine-tunes models to produce outputs consistent with user expectations (Ouyang et al., 2022). However, alignment doesn’t end at training. Studies show that users often push back on AI responses that conflict with their values, steering behavior through real-time feedback, a process termed user-driven value alignment (Fan et al., 2025). Such interactive alignment is critical in high-stakes contexts like medical decision support, where LLM-powered co-pilot systems have helped clinicians catch more errors than either humans or AI alone (Ong et al., 2025). These findings show the importance of transparency, oversight, and human agency in AI deployment.
Another approach, Constitutional AI, replaces constant human feedback with rule-based self-supervision. Models are trained to revise outputs using explicit principles like “Be respectful” (Bai et al., 2022). Recent work has expanded this into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024). Tools like ConstitutionMaker further help users turn natural language feedback into usable policies (Petridis et al., 2024). DoubleAgents contributes to this space by embedding alignment into interactive simulation. Users iteratively refine coordination policies, email tone, and stop hooks through rehearsed ReAct loops, treating alignment not as a one-time step, but as a continuous, participatory process.
2.2. Agentic AI in Task Planning and Execution
AI agents, systems that perform tasks on users’ behalf, operate through tool use (e.g., APIs) (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). While these agents show growing capabilities, delegating control to them still raises concerns around trust, transparency, and alignment (Afroogh et al., 2024). Planning and execution remain core challenges. Benchmarks like Tau-Bench, NaturalPlan, TravelPlanner, and AppWorld (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024) highlight the difficulty of robust multi-step planning, especially in open-ended tasks. To address this, methods like task decomposition (Zhang et al., 2024) and ReAct, which interleaves reasoning and acting, have been shown to improve agent adaptability (Yao et al., 2023), especially when used in collaborative or multi-agent domains (Lal et al., 2023).
However, fully autonomous agents like AutoGPT often struggle with robustness and transparency in real-world use (Zheng et al., 2024; Yao et al., 2024). Cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental and context-sensitive planning. Recent studies show that exposing agent reasoning and allowing user intervention improves both performance and trust (He et al., 2025). In sum, planning in agentic AI is not just a technical hurdle; it is a design problem. Transparent reasoning, iterative refinement, and human-in-the-loop oversight are essential for building trustworthy agents that can meaningfully support real-world tasks.
2.3. Human Simulations with LLMs
Recent work has shown that LLMs in multi-agent simulations can exhibit surprisingly human-like behaviors, forming routines, organizing events, and coordinating spontaneously without explicit instructions (Park et al., 2023; Horton, 2023; Park et al., 2022). These behaviors often reflect human reasoning and social strategies, aligning with patterns observed in economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial decisions (Hamilton, 2023), and over 80% of psychology experiments (Hewitt et al., 2024). This suggests LLM agents can serve as useful proxies for modeling human behavior at scale.
Beyond simulation, LLMs are demonstrating real-time social competence. CICERO, a dialogue-enabled Diplomacy agent, ranked in the top 10% of human players by integrating language with strategic planning (Team et al., 2022). Similarly, the MATRIX framework used simulated AI dialogues to surface moral tradeoffs in Constitutional AI evaluations (Pang et al., 2024), showing how LLMs can assist in value alignment. LLM simulations also support red teaming and adversarial testing, enabling safe, scalable evaluation of edge cases and vulnerabilities (Purpura et al., 2025). While imperfect, these agents provide a reproducible and low-risk method for probing system behavior in socially sensitive contexts, offering a powerful complement to human-centered alignment and evaluation.
3. DoubleAgents System
Figure 3. A screenshot of DoubleAgents coordinating the assignment of four speakers to four seminar slots. (A) Policy Panel – Displays coordination policies that reflect user values and preferences. Users can add, edit, or delete policies, which guide the reasoning and planning of the coordination agent. (B) Interactive Chat Interface – Shows the full history of plans and actions, allowing users to monitor planning and execution, confirm or modify agent suggestions, and intervene when necessary. (C) Plan and Action Pop-up – A focused view within the chat history where users can review, edit, regenerate, or approve individual plans and actions. (D) Assignment Tracker – Visualizes tentative speaker-to-slot assignments and indicates each speaker’s availability. Dark purple denotes confirmed availability; light purple signals ambiguous availability requiring user confirmation. (E) Communication History Viewer – Displays the messaging history between the coordination agent and each speaker, organized by date and interaction. (F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks. In this paper, we focus on a concrete use case: assigning seminar speakers to available time slots (see Appendix A for additional background). DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies. Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows. The system comprises six core components organized into three functional categories. First, DoubleAgents includes the Core ReAct Agents, which implement the reasoning, planning, and simulation logic that drives the coordination process:
  • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
  • Simulated Respondents - Represent background agents that emulate human respondents’ behavior by responding to user-generated emails based on predefined personas and descriptions.
Second, DoubleAgents incorporates the System Orchestration Modules, which manage the internal context and ensure robustness during decision-making:
  • Context Management Module - Maintains all relevant contextual information related to the task, including simulated respondents’ details (e.g., personas and descriptions), email logs from prior time steps, and the history of user-approved plans, actions, and emails.
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions. These escalations are recorded as stop hooks, enabling iterative refinement and extension of the policy set.
Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can approve or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
  • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.
In this section, we provide a walkthrough example to illustrate how DoubleAgents operates in practice. We also describe the implementation details of DoubleAgents. Consider the example of Lee, a PhD student organizing the HCI seminar series at his university. Lee plans to invite four speakers–Prof. A (a famous professor), Prof. B (a rising star professor), Student X (a senior PhD student), and Student Y (a PhD student who is a friend of Lee) – to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
Lee begins at the landing page, where he specifies his goals to the system. He types: “I am trying to organize a seminar with four speakers and four slots, help me reach out”. He then provides additional details, including the seminar location, and the dates of the four time slots. Next, Lee enters persona information for each speaker he wants to invite, which the system uses to generate simulated speaker profiles. The landing page is shown in Figure 4. To further customize communication, Lee uploads five of his past emails, allowing the system to adapt its generated messages to his style; for instance, casual for friends and formal for faculty. Finally, Lee initiates the simulation by clicking the “Start Planning” button, prompting the system to set up the scenario with the specified speakers and time slots. At this point, he is also shown a list of preset policies (see Figure 3(A)), which serve as guidelines for the coordination agent to begin planning. Lee reviews the policies, all of which appear reasonable. He may modify existing policies or introduce new ones to better guide the agent’s planning process.
Figure 4. The landing page to specify the user goals, seminar details, and the persona info for the speakers.
3.1. Context Management Module
At the beginning of the scheduling process, all information provided by Lee on the landing page is compiled into a structured problem context that the system consults at each step to solve the assignment task. DoubleAgents represents this problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history, etc.), and each value contains the corresponding content. As shown in Figure 2, an example snapshot of the problem context includes keys such as ”time info”, which reflects the seminar slots specified by Lee. The context is dynamically updated throughout the interaction, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
3.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 3(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 3(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix C.4.
(a) The pop-up window for a generated plan.
(b) The pop-up window for generated actions following the plan.
(c) The pop-up window for a generated email.
Figure 5. Examples for plan and action generation.
3.3. Coordination Agent
3.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix C.1.1.
3.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, the lead authors brainstorm and define rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 3(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix C.1.2.
3.3.3. Plan Generation
Guided by this policy, the coordination agent generates the following plan (shown in Figure 5(a)) for Lee: “Send email to Prof. A, Prof. B, Student X, and Student Y to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”. Lee considers the plan and the rationale logical, since scheduling cannot proceed without first collecting availability information, and clicks ’Approve & Continue’ to move forward. The plan and its rationale are automatically recorded in the interactive chat interface (shown in Figure 2(B)), allowing Lee to review past activities and understand the rationale behind the plan. If Lee is dissatisfied with the proposed plan, he can either request the system to regenerate an alternative or directly edit the current plan himself, as shown in Figure 5(a).
3.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 5(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix C.1.3.
3.3.5. Action Execution
Following approval, the coordination agent proceeds to instantiate each action by generating email drafts for the designated speakers, including both subject lines and message bodies (Figure 5(c)). This is done by calling the email writing tool, as specified in the tool box shown in Figure 2. Specifically, it calls the writeEmail() API, which drafts emails given the current problem state, seminar details, user-provided templates, and historical email edits. These elements allow the system to learn and adapt to the Lee’s preferred communication style, capturing tone, level of formality, and domain-specific phrasing, when generating new emails. Note that email drafting is performed using GPT‑4o, which is invoked by the coordination agent during the execution of the writeEmail() tool.
Lee reviews each draft and identifies one instance where the tone is not appropriate—the email to Prof. A, a distinguished professor, feels overly informal. To address this, Lee types: “Please make the email sound more formal, as Prof. A is a distinguished professor I’d like to invite.” The system then regenerates a more suitable version. Lee further refines the message by editing the first paragraph to better reflect his personal voice: “My name is Lee, a PhD student in Computer Science. Your work in [XXX] has greatly inspired us, and I would be honored to invite you to speak at our HCI seminar series.” After reviewing and finalizing all drafts, Lee is satisfied and clicks ‘Approve & Send’, prompting the system to dispatch the emails to all invited speakers. This is accomplished by invoking the sendEmail() API by the coordination agent, targeting the appropriate speaker email addresses specified in the problem context. In addition, all email-related activities are logged in the dashboard (Figure 2(E)), enabling Lee to easily monitor the status and history of interactions with each speaker. Besides writing and sending emails, the coordination agent may invoke actions involving delayed responses, such as waiting for a speaker’s reply. This is achieved by calling the waitResponse() tool. In such cases, no message is generated or sent; instead, the agent pauses execution and either proceeds to the next pending action or advances to the next simulated day. Overall, we use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
3.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix C.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
3.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 2), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix C.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 6(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
(a) Email sent by the simulated Prof. A.
(b) Edge case detection given response from Prof. B asking about Zoom options.
Figure 6. The example email sent by the simulated respondents and the system message when detecting edge case.
3.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. Lee’s clarification is recorded as a stop hook and used to refine the policy set, explicitly specifying that Zoom is not preferred when speakers request remote participation. In this way, user interventions during edge cases are converted into reusable system policies. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix C.2).
3.6. Dashboard Visualization
To support user understanding and foster alignment, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 3(D), (E), and (F), which we describe in detail below:
  • Assignment Grid and Progress Tracker (Figure 3D): This panel visualizes tentative and confirmed speaker-to-slot assignments. It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled. This layer helps Lee assess how close the current plan is to completion at a glance.
  • Calendar and Timelines (Figure 3F): Anchored in simulated days, this temporal layer helps Lee reason about timing, deadlines, and follow-up actions. The calendar highlights both the current simulated day and the seminar dates, allowing Lee to prioritize actions based on urgency and scheduling constraints.
  • Communication Visualization (Figure 3E): This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status. It enables Lee to track ongoing communications, detect gaps, and assess follow-up needs.
These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time. Throughout the process, Lee interacts with the coordination agent and the respondent agents to collaboratively establish speaker-to-slot assignments for the seminar. Together, these visualization features provide Lee with a clear overview of assignment progress and a sense of what needs to be done on the current and subsequent days.
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix C.5. As shown in Figure 3(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
4. Technical Evaluation
DoubleAgents enables users to engage in iterative task planning and action execution through two core functionalities: (1) policy selection, which guides the generation of plans aligned with predefined coordination strategies; and (2) robust edge case detection, which identifies problematic or ambiguous situations and escalates them to the user for clarification and resolution. In this section, we describe the technical evaluation of the two components.
4.1. Policy Selection Experiments
4.1.1. Experiment Settings
Policy selection refers to the process of identifying relevant coordination policies that guide the generation of high-quality, context-appropriate plans. Specifically, given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module is expected to retrieve all applicable policies that can serve as guidelines for the next planning step. For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts. To inform policy selection, we use the output of the getProgressSummary() API, which provides a high-level snapshot of the current task state. An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07. All other speakers are still pending. Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix D. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.
4.1.2. Results and Findings
As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state. The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection. Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval. Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
Prompting Techniques F1 Score
Mean SD
Speaker-only .27 .054
Slot-only .11 .038
Slot + Speaker .27 .017
Summary-based .70*** .035
Table 1. The technical evaluation results for Policy Section comparing four prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
Prompting Techniques Accuracy
Mean SD
Zero-shot .79 .021
Few-shot w/o Policy .88† .000
Few-shot w Policy .93** .015
Table 2. The technical evaluation results for Edge Case Detection comparing different prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
4.2. Edge Case Detection Experiments
4.2.1. Experiment Settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix D.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix D.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
4.2.2. Results and Findings
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Together, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
5. User Study
5.1. Goal and Overview
To investigate how users engaged with DoubleAgents’ features and how interactive simulation shaped their use of agentic behavior, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their understanding and reliance on the system, and how their understanding and calibration of agent behavior evolved over time. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences via their policies, email templates, and stop hooks. Participants completed two questionnaires. First, an AI comfortableness survey measured trust and alignment needs for proactive actions and comfort with automating planning and execution, administered before Session 1, after Session 1, and after Session 2. Second, a System perception survey captured UX and perceptions of 15 system features after Session 1 and after Session 2, enabling us to track how repeated use shifted alignment strategies, delegation boundaries, and reliance. Detailed study procedures and data analysis methods are provided in Appendix A.
5.2. Participants
We recruited 10 participants (7 female, 3 male; aged 21–30, M = 25.7) from our university community. All participants (10 of 10) reported using generative AI automation tools daily. Participants primarily used tools such as AI writing assistants (10 of 10), AI coding assistants (7 of 10), and AI design/creative tools (5 of 10). All participants reported solid understanding of generative AI: 7 described having in-depth knowledge of models and tools, while 3 reported a good general understanding of key concepts. Participants’ experience with human coordination tasks varied widely. 2 participants had moderate to extensive experience, having led team schedules or group planning multiple times. 3 participants had some experience, having organized small tasks or events independently. 4 participants had minimal experience, primarily helping or shadowing others in coordination tasks. Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
5.3. Result
5.3.1. Consistently Positive User Experience
Across both sessions, all 10 participants reported a consistently positive experience with DoubleAgents. Overall system ratings remained high and showed a slight improvement (from 6.30/7 to 6.62/7). Participants also perceived the system as effective in generating optimal plans, actions, and emails (5.90/7 to 5.88/7), and judged it as robust (5.60/7 to 5.88/7). Both mental and physical demand remained low across sessions, decreasing from 2.90/7 to 2.13/7 and from 1.50/7 to 1.13/7, respectively. Participants reported high perceived performance (6.30/7 to 6.75/7), low frustration (1.80/7 to 1.25/7), and a near-significant reduction in effort (3.20/7 to 1.88/7, $p=0.054^{{}^{\wedge}}$). Together, these results indicate a consistently positive and efficient user experience across sessions, with effort showing the strongest improvement.
5.3.2. Increasing User Reliance Driven by DoubleAgents’ Design
Despite largely stable system-level scores, participants’ perceived reliance on DoubleAgents increased significantly over sessions (from 4.90/7 to 5.75/7, $p=0.02^{*}$). Participants characterized this shift not as “unconditional reliance”, but as increased comfort with delegation as the system’s behavior became more user-aligned, legible, and shapeable—enabled by clearer controls, predictable execution patterns, and explicit points of intervention. A correlation analysis between reliance and feature-level perceptions (Spearman; FDR-corrected) showed that reliance was associated with features supporting control and explainability, including state visualization ($r=0.78$, $p<0.05^{*}$) and confirmation-based user control ($r=0.51$, $p<0.05^{*}$). For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker. She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.” Increases in reliance over time were strongly linked to DoubleAgents’ human-AI alignment features. In particular, reliance increased with perceived system robustness demonstrated through interactive simulation ($r=0.54$, $p<0.05^{*}$), and was reinforced by repeated successful issue flagging via stop hooks ($r=0.65$, $p<0.01^{**}$) and perceived value alignment through coordination policies and email writing ($r=0.54$, $p<0.05^{*}$). Reliance was negatively correlated with mental demand (r=0.47r=-0.47, $p<0.05^{*}$) and showed no significant correlation with other system experience or perception items. Collectively, these findings suggest that reliance increased not merely due to perceived system capability or general UX, but because of DoubleAgents’ support for alignment in design. The following section details how participants experienced and leveraged this alignment through interactive simulation.
Figure 7. Average user perception scores toward comfort with automating the (a) planning and (b) execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$, * at $p<0.05$, ** at $p<0.01$, *** at $p<0.005$.
5.3.3. Increasing User Comfort with AI Automating Human Coordination Tasks: Contributing Alignment Supports.
Over interactive simulations, participants became more comfortable delegating both planning and execution tasks to agentic AI. For planning (Figures 7a, participants became more comfortable letting the AI visualize current status (4.80/7 to 6.50/7, $p<0.05^{*}$), propose time assignments (5.10/7 to 6.38/7, $p<0.1^{\vee}$), decide daily next steps (4.70/7 to 6.50/7, $p<0.005^{***}$), and monitor/flag issues (5.50/7 to 6.50/7, $p<0.01^{**}$). For execution, comfort increased for drafting-only emails across all four types (from around 6.0 before to nearly 7.0/7 after, $p<0.01^{\vee}$), and also increased for draft-and-send tasks on users’ behalf—e.g., initial outreach rose from around 2.7/7 to 4.6/7 ($p<0.05^{*}$), with similar gains for follow-ups and final confirmation ($p<0.1^{\vee}$).
Participants attributed this increased comfort to DoubleAgents’ alignment supports: interactive simulation; user-editable coordination policies that guide next-step decisions; email templates and style controls that realize those actions; and stop hooks for issue escalation. Together, these supports advanced bidirectional alignment: the AI agent’s actions became better aligned with users’ needs and values, and users developed greater understanding and trust, becoming more comfortable with the AI taking initiative in automating this human coordination task.
Interactive simulation as a testbed for policy, email, and stop-hook calibration. Simulations supported alignment by letting participants safely probe edge cases, identify where the agent’s behavior diverged from expectations, and iteratively refine (1) what the agent should do (coordination policies), (2) how it should carry out those actions in email (templates and style), and (3) when it should pause for confirmation (stop hooks). Participants emphasized that interactive simulations increased confidence in their perception of the system’s robustness (5.80/7 to 6.38/7). They found simulated test cases realistic and representative of real-life scenarios (5.90/7 to 6.75/7), and especially useful for surfacing potential problems (5.80/7 to 6.50/7). P3 described how simulation exposed realistic exceptions and prompted advance policy-setting: “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule—it forced me to think about how I would handle that in real life and to create relevant policies, and that was good. I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
Email templates and coordination policies fostered increased comfort with AI automation. Participants’ comfort with AI automating planning and execution increased over time as the system’s actions became more consistent, better grounded in the users’ own values and rationale, and increasingly “in their voice.” Carrying over value-aligned coordination policies made participants more comfortable delegating planning—e.g., having the AI decide daily next steps, track and visualize progress, make assignment decisions, and monitor/escalate issues when cases did not fit existing policies (Figure 7a). Email-template carryover particularly supported execution delegation. Comfort increased for drafting-only emails and for draft-and-send tasks (Figure 7b). One exception was drafting and sending clarification emails without any user confirmation, which showed little change (3.00/7 to 3.90/7 to 3.90/7). P10 noted that clarifications often require nuanced, user-specific judgment: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” As a result, participants preferred to remain in the loop for clarification messages. Overall, participants described an early sense of being understood and represented by the system, as if it “got them,” when iterating on email templates and coordination policies. For example, P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” Similarly, after refining policies, P3 observed that the agent’s choices matched their expectations: “I found that the agent’s recent actions followed exactly the policies I had in mind—super aligned.” This growing alignment helped the system reflect their intended actions and communication style, increasing their comfort with offloading tasks to AI agents.
Stop hooks helped establish safe delegation boundaries, fostering users’ comfort with delegation. Participants highlighted stop-hook flagging as particularly important for maintaining safe delegation boundaries. For monitoring and surfacing potential issues, P1 described issue flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress—when they skimmed the respondent agent’s reply and thought everything was resolved, the flag showed that it was not. Thus, with these “stop hooks” continuously monitoring for exceptions and prompting intervention when needed, participants felt much more comfortable offloading tasks to the AI.
Figure 8. Average user perceptions of alignment needs, expectations, and trust in agentic AI systems (average scores across 10 users: before Session 1, after Session 1, and after Session 2). * denotes statistical significance at $p<0.05$ and ** at $p<0.01$.
Increasing Expectations and Needs for Alignment in Agentic AI
Across sessions, participants reported stronger expectations for how proactive AI in general should behave and align with them (Figure 8). In the AI Comfortableness Survey, participants reported a significant increase in wanting the AI’s actions to align with their preferences, values, and priorities, from 6.10/7 (before using the system, based on their previous AI experiences) to 7.00/7 after two uses of DoubleAgents ($p<0.05^{*}$). Participants’ desire to stay in control remained high (5.30 to 5.75). They also reported an increased need for detailed system rationales and explanations before the system acts, from 5.00/7 to 6.00/7 ($p<0.05^{*}$). Participants framed clear and abundant explanations as a core alignment mechanism for building an accurate mental model of the AI’s decision process, enabling verification, and determining when and how to intervene. P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency… if I didn’t get that, I would feel lost [and] not know what’s going on…” We interpret this as evidence that DoubleAgents’ interaction design helped participants recognize the importance of alignment and articulate stronger alignment expectations for future proactive AI usage.
Furthermore, participants also reported statistically significant increases in trust to let similar AI systems act on their behalf, from 3.20/7 to 5.00/7 ($p<0.01^{**}$), even as their assessments of AI reliability remained relatively the same. This suggests that alignment features can increase comfort with delegation beyond improvements in perceived raw capability. P1 noted that even though DoubleAgents represented only one specific experience, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.”.
6. Deployment Studies
6.1. Goal
As the in-lab study demonstrated, the simulation was helpful, and participants felt they were getting more comfortable use agentic AI help planning and executing human coordination tasks. However, would real-world seminar organizers find DoubleAgents useful and relevant? If so, how would they use it-would they plan, draft, and send invitation, follow-up, or clarification emails in different ways? Would alignment be established in a similar manner? What additional challenges might arise in real-world use? To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited organizers who were managing actual events across three real-world deployments.
6.2. Deployment Participants and Procedures
6.2.1. Use Cases and Participants
In our deployment study, DoubleAgents was used in three real-world event-planning workflows involving two organizers (P11, P12). P11 used DoubleAgents to coordinate a Fall 2025 HCI seminar series (Aug 2025; Deployment 1, D1), inviting 10 speakers from a pool of 23 to fill 10 slots; we observed and logged system use for 7 days during this process. P11 later used DoubleAgents again for a Spring 2026 HCI seminar series (Jan 2026; Deployment 3, D3), inviting 11 speakers from a pool of 20 to fill 11 slots; we tracked use for 11 days. P12 used DoubleAgents to organize a conference workshop panel (Sep 2025; Deployment 2, D2), inviting 5–6 panelists from a pool of 18 to fill a single slot; we tracked use for 7 days. We recruited P11 and P12, two experienced seminar organizers (one male, one female; average age 31), via personal connections. Both were PhD students who had previously organized departmental research seminars and had prior experience using generative AI for writing emails.
6.2.2. Introduction, Setup, and Interactive Simulation Rounds
Organizers were first informed about the study and provided consent. They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops, illustrated with a simulation example from the user study. Organizers also completed the initial survey covering AI perception and demographics.
To set up the system, we disabled the respondent module from the lab version of DoubleAgents and connected the SendEmail() and checkEmailUpdates() tools to Gmail’s SMTP service, enabling organizers to send emails and receive replies through their personal Gmail accounts.
For the simulation rounds, organizers prepared a simulation spreadsheet using their real organizing data (including real potential speakers’ names), first with placeholder email addresses and later with the real ones, and provided their Gmail SMTP credentials. Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw. Before connecting their real email accounts, participants completed a 2-hour interactive simulation session using pre-generated speaker personas based on the real candidates, allowing them to safely familiarize and themselves align with the system’s workflows. Finally, they tested the real email-sending module twice by sending messages to placeholder addresses (the researchers’ and organizers’ personal emails) to confirm successful delivery; afterward, we switched the recipient addresses to the invitees’ real Gmail accounts.
6.2.3. Live Email Execution and Response Handling
After the simulations, organizers used DoubleAgents to initiate the coordination process and track it for over a few days. They continued receiving system updates over the following week (D1: 7 days, D2: 4 days, D3: 11 days). We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning. Organizers’ system interactions, including email sends, scheduling updates, and system adaptations, were tracked. Qualitative feedback was collected via check-in meetings, while quantitative system logs were collected for system learning over the week.
6.3. Deployment Results
6.3.1. DoubleAgents’ Perceived Usefulness and Relevance in Real-Life Human Coordination Problems.
After the interactive simulations, both organizers (P11 and P12) reported that DoubleAgents was highly relevant to their existing seminar-organizing workflows. They highlighted that the system fit seamlessly into their email-heavy tasks, and that its control settings and edge-case flagging helped them handle complex coordination scenarios that conventional scheduling tools (e.g., Calendly) cannot easily manage. P11 emphasized that crafting good email templates is a delicate and important task because it requires personalizing messages to “make the speakers feel they are special,” even when requesting a favor. Both P11 and P12 found DoubleAgents capable of generating strong starter emails for them to edit early on, as well as producing acceptable final emails to send. P12 especially appreciated having explicit coordination policies in the system to guide next-step decisions in ways that matched his decision-making habits and values. He also found it useful that rationales were always visible and editable, in contrast to AI tools that hide their reasoning. After using DoubleAgents, P12 remarked, “I trust AI [systems] for specific use cases more,” explaining that he found DoubleAgents dedicated to and capable of generating high-quality emails. P11 found the color-coded visualization of speaker constraints and response states (e.g., not yet outreached, tentative yes, yes, no) especially useful; after D1, she adopted a similar color-coded scheme for other tracking tasks she previously managed in spreadsheets.
6.3.2. Interactive Simulation as an Alignment Method for Complex, Realistic Tasks.
Both P11 and P12 emphasized that interactive simulation was important for understanding DoubleAgents and for setting up initial coordination policies and email templates. They appreciated that simulations used real invitee names and incorporated persona cues and edge cases (e.g., funding requests, cancellations, Zoom requests), which helped them anticipate challenges and craft templates that would transfer to live deployment. For example, P12 used early simulation rounds to iterate on what information to include and where to place key details (e.g., the workshop webpage link) in his email templates, similar to how he had previously tested invitation templates by emailing himself.
P11’s D1 simulation surfaced multiple policy- and tone-sensitive situations (e.g., rescheduling after a slot had already been assigned), which prompted iterative template and policy revisions. P11 reported that the system’s drafts were largely aligned with her language, but she removed or adjusted certain lines to better match her preferences and to clarify what information should be included when edge cases occur. Overall, P11’s D1 interactive simulation took four days across five meetings: the first two meetings focused on ensuring the system ran smoothly for a high-complexity scenario (20+ candidates, 10 slots, and substantial context per speaker). During setup, P11 revised invitation and follow-up templates more than four times and updated speaker descriptions so email generation could incorporate richer context. After we refined the system and onboarding protocols, P12’s D2 simulation took two days across two meetings. P12’s panel invitation task was also simpler (e.g., shorter descriptions and a single-slot panel), which further reduced the needed time. Both organizers reported that practice simulation rounds were valuable, and that sending test emails via SMTP to his own personal gmail helped him gradually understand system capabilities and limitations. Overall, these deployments suggest that simulation and onboarding effort scales with task complexity and the amount of required context.
During D1, P11 specifically highlighted the importance and difficulty of providing sufficient context for highly personalized emails, especially for real-world cases. She noted that “[persona description management] is really tricky.” For instance, one invitee had previously coordinated with a different organizer in a prior semester (Friday availability), but the D1 seminar series had shifted to Thursdays. To avoid missing or conflicting context across past threads, schedule changes, and templates, P11 updated the persona description used for simulation with an additional line noting that the invitee had been told the wrong day in the past (Friday 4–5pm) and that the seminar was now on Thursday 4–5pm. In another case, P11 had already discussed the seminar with a speaker on Discord but initially omitted this context during early simulations; she then added it to the persona description so the follow-up email could adopt the appropriate tone and content. After rehearsing with real-speaker data in simulation, P11 finalized a preferred template for live sending.
P11 also noted that even in D3 (her second deployment), simulation remained necessary because “real life is super complex.” She observed new patterns in edge cases, such as delayed responses or vague “check back next month” replies from senior faculty, which led her to revise policies (e.g., longer follow-up intervals) and adopt a more formal email tone. Across both deployments, simulation served not just as a testing ground, but as a means of progressive alignment: helping users adapt policies, templates, and expectations to shifting social and contextual dynamics. These findings suggest that onboarding and setup effort scale with task complexity and that simulation continues to play an ongoing role in supporting user-aligned adaptation, even after initial deployment.
6.3.3. Example: Live Email Execution, Response Handling, and Edge Cases
During D3, P11 sent emails to 11 selected speakers and relied more heavily on templates and system-generated drafts on Day 1 (a Friday), reflecting increased familiarity with the workflow. DoubleAgents then suggested waiting for responses over the following two days due to the weekend, which P11 accepted. Late on Day 3 (Sun), Invitee J and Invitee A replied with questions, which DoubleAgents surfaced in P11’s Day 4 (Mon) DoubleAgents session. On the morning of Day 4, Invitee B replied with availability. Based on these responses, DoubleAgents suggested replying first to Invitees A and J, but not to Invitee B, to preserve future scheduling flexibility and avoid potential conflicts. P11 accepted this suggestion and drafted replies accordingly. Specifically, Invitee A asked a funding-related question and included a personal greeting (e.g., “Happy New Year… if it doesn’t materialize…”). P11 revised the AI-generated reply by incorporating similar phrasing from their own writing style, but chose not to save this edit as a reusable template, noting that it reflected a personal, context-specific language choice. On Day 5, Invitees I and J replied, and DoubleAgents suggested allocating them to earlier-semester slots, which P11 accepted and responded to. Similar reply-and-assign cycles occurred on Day 6. On Day 7, DoubleAgents suggested following up with all the invitees who had not yet responded, in accordance with P11’s newly updated D3 policy to follow up every four working days (rather than every two, as in D1). P11 proceeded using a personalized follow-up template that incorporated questions added on Day 4. Also on Day 7, P11 handled additional questions and confirmed slots with Invitees C and H, who had replied promptly with apologies and availability. Later that evening, Invitee I emailed to report a schedule change and requested reassignment to an unclaimed date. On Day 8, DoubleAgents suggested reassigning Invitee I and requesting additional availability from Invitees H and B to resolve emerging conflicts. P11 accepted and carried out these actions. Over the weekend (Days 9 and 10), Invitee H responded with new dates but noted that P11 should follow up again the following month, as their schedule was still tentative. DoubleAgents accordingly scheduled a follow-up for Day 35, while on Day 11 it prompted P11 to address remaining unresponsive and unresolved cases.
Collectively, these interactions illustrate the complexity and multi-threaded nature of this real-world task. As P11 reflected, “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful. The new policies and templates I developed through this semester’s simulation—since everyone is a professor—worked really well. The simulation part cannot be skipped.” This case highlights DoubleAgents’ strong use case, demonstrating how simulation, policies, and email templates work together to support alignment between the system’s actions and users’ intentions.
7. Discussion
7.1. Human-AI Alignment via Interactive Simulations
As AI systems evolve from passive tools to agentic collaborators that plan and act on users’ behalf, human–AI alignment becomes not just a technical challenge but an interactive one. Our findings suggest that interactive simulation offers a compelling design mechanism to support this alignment by foregrounding user agency, surfacing system behavior early, and fostering a co-adaptive relationship between humans and AI.
Across our lab study and deployments, we observed that interactive simulation enabled users to engage with alignment as a gradual process, rather than a one-time configuration step. Participants entered the system unsure of their own preferences and reluctant to delegate; yet through repeated simulated cycles, each revealing system actions, uncertainties, and edge cases, they clarified values, authored and refined policies, and ultimately built enough user comfort to let the system act on their behalf. This shift mirrors real-world dynamics of human delegation: people do not fully trust a new assistant immediately but gain confidence through concrete interactions over time.
Simulation thus serves multiple alignment functions simultaneously. First, it externalizes system behavior, making invisible decision logic (e.g., why the agent followed up or chose a certain speaker) legible to users. Second, it creates safe interaction spaces, akin to usability testing for AI plans. Like sandboxed environments in traditional software testing, simulations help users anticipate downstream consequences without real-world stakes, crucial in settings like email coordination where tone, timing, and etiquette are socially sensitive. Third, simulation supports user-side alignment, helping users identify preferences they hadn’t articulated upfront (e.g., when to prioritize seniority, how long to wait before following up, or what tone feels appropriate across contexts). Rather than requiring exhaustive foresight, DoubleAgents invites users to iterate on policies as new situations emerge.
This design reframes alignment as a human-AI collaboration process: the agent proposes actions, but the human remains in the loop, interpreting, approving, and updating the future behavior of the system. In contrast to approaches that aim to extract user goals from demonstrations or feedback, we emphasize the value of direct, interactive co-construction of alignment artifacts, such as stop hooks, policy rules, and email templates, which evolve with use cases. This opens new design directions: how can we better scaffold early-stage alignment for users who don’t yet know their own preferences? What kinds of interfaces help users develop trust while retaining oversight? How might simulation be extended to other high-stakes or interpersonal agentic domains, such as hiring, negotiation, or caregiving? Our work suggests that alignment is not a prerequisite to agentic AI, but a continuous, situated interaction practice. By embedding simulation into the human–AI workflow, we give users the tools not just to observe and react to agent behavior, but to actively shape and steer it. In doing so, DoubleAgents contributes to a broader vision of agentic AI that is transparent, participatory, and grounded in human values from the start.
7.2. Designing User Onboarding Experiences for Agentic AI
Onboarding is a critical site for agentic AI systems. Unlike traditional tools, agentic systems require users to make early decisions about goals, decision boundaries, tone and formality, and what to delegate. They also create distinct discoverability challenges: users often do not know what the agents can do and what they can do. Key modes and safeguards (e.g., auto-send vs. draft-only) may be implicit, and failures can appear only in rare or edge cases—so there is no single, linear “progress bar” that reveals system status or readiness. As a result, effective onboarding is necessary for users to form accurate mental models, judge reliability, and feel comfortable delegating and deploying the agent in higher-stakes settings (e.g., a seminar organizer connecting to Gmail to send real invitations to senior faculty—who may also be potential postdoc advisors—where mistakes can harm professional relationships and reputation). In DoubleAgents, we found that interactive simulation provides a powerful onboarding scaffold, helping users observe, evaluate, and refine system behavior before real-world use.
Rather than presenting configuration as a static setup phase, DoubleAgents embeds onboarding through repeated rounds of interactive simulation. During these rounds, edge cases (e.g., a speaker canceling or requesting Zoom) expose users to socially complex scenarios, help them become familiar with the system’s behavior, and elicit their personal preferences. Participants in our study often hesitated to delegate initially, not due to lack of interest, but due to uncertainty about agent behavior. Using repeated rounds of interactive simulation as an onboarding task lowered this barrier by helping participants move past the initial learning curve and novelty effects, and by revealing how the agent behaved across diverse cases. Like training a new assistant, onboarding becomes a co-configuration process in which expectations, autonomy, perceived usefulness, and trust are developed together.
We argue that onboarding for agentic AI should not be treated as a brief tutorial animation or tooltip overlay. Instead, it can be designed as a structured, interactive process where users and systems collaboratively define goals, boundaries, and escalation rules. DoubleAgents shows how early simulations can accelerate alignment, ease delegation, and set the foundation for long-term, trustworthy collaboration.
8. Limitations and Future Work
Our work has several limitations that point to promising directions for future research. First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios. While these simulations were reviewed by domain experts and designed to reflect realistic coordination challenges, they cannot fully capture the diversity and unpredictability of real-world interactions. Future work could strengthen alignment by grounding both the coordination agent and respondent simulations in users’ real historical data, such as past email threads, organizing decisions, and communication patterns. This would enable more personalized modeling of users’ values, tone, and escalation preferences, as well as more behaviorally faithful respondent agents. Second, DoubleAgents currently relies on different language models for different subtasks (e.g., o3 for reasoning and 4o-mini for generation). Although our technical evaluation suggests this division performs adequately, future models with stronger social reasoning and long-horizon planning capabilities may further improve coordination quality. In addition, tighter integration between user-specific data and model reasoning, such as learning when and how to apply policies from prior interactions, remains an open challenge. Third, while we conducted real-world deployments with two seminar organizers, discussions with additional organizers revealed coordination scenarios that our current system does not fully support. For example, bilingual invitations, multi-organizer decision-making, and negotiation among organizers introduce shared authority, language switching, and conflicting preferences. Extending DoubleAgents to support collaborative alignment, where policies, stop hooks, and responsibilities are jointly defined and negotiated, represents an important direction for future work. Finally, all study participants were experienced users of AI tools, which may have positively influenced their willingness to delegate and their trust in the system. Future studies should include participants with a wider range of AI familiarity to better understand onboarding challenges, mental model formation, and how interactive simulation can support alignment for users with less technical background.
9. Conclusion
We introduced DoubleAgents, a system for aligning agentic AI through interactive simulation, policy-grounded planning, and transparent oversight. In coordination tasks like speaker scheduling, alignment is not just about correctness; it’s about respecting user preferences, adapting to edge cases, and maintaining trust. By embedding simulation into the agent workflow, DoubleAgents helps users rehearse edge cases, refine their preferences, and gradually scale up delegation. Our lab and real-world deployments show that users increasingly trusted and customized the system, not because it was perfect, but because they could iteratively shape it. This work offers a practical path toward alignment in real-world agentic AI: one grounded not in static prompts, but in interaction, iteration, and co-adaptation.
10. Generative AI Usage Disclosure
We used generative AI tools (e.g., OpenAI’s ChatGPT) as support during several parts of this project. Specifically, we used them to help debug the system, draft non-critical implementation and UI code, write analysis scripts for logs and study data, and edit text for grammar and clarity during paper writing.
Appendix A Seminar Planning: A Human Coordination Problem
Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate. Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time. Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—derail coordination and strain relationships. These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
We study these challenges through one concrete and consequential setting: organizing an academic seminar series. A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term. Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own preferences-such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
However, the most unpredictable and challenging part is dealing with responses. Some people do not reply at all and require multiple follow-ups to elicit a response. Some give only vague and tentative replies that need repeated confirmation. Some can be flakey and unreliable, offering slots but later canceling. And many unexpected problem cases arise: people might be on parental leave, ask for visa letters or funding support, recommend a different speaker, request a different date or talk modality, or raise logistical demands. In such human coordination tasks, you can rarely predict what is coming. Therefore, conventional scheduling tools such as Calendly [^1] https://calendly.com/ or When2Meet [^2] https://www.when2meet.com/ are ill-suited to this setting. They assume stable constraints and one-shot coordination, yet seminar scheduling requires ongoing negotiation, persuasion, and relationship maintenance. Unlike ordinary meeting scheduling, seminar speakers are doing you a favor, which means you cannot simply send them a link to pick a time—you must persuade and negotiate. Social dynamics and etiquette also play a large role: organizers must time their emails carefully to avoid overwhelming, annoying, or losing momentum with potential speakers, who are often busy scholars. Knowing when to nudge, follow up, or wait is subtle. Email remains the standard medium for professional and academic outreach, allowing organizers to adapt tone—formal, casual, or enthusiastic—to different speakers in ways scheduling links cannot. It also supports sharing rich context, such as talk topics, background, past speakers, or travel plans—far beyond simply choosing a calendar slot.
Thus, experienced organizers adopt personal strategies to cope with such a demanding task. For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again. Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints. When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption. For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability. Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate. An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
Appendix B User Study Procedures and Data Analysis
Test Scenario Preparation.
We selected eight distinct personas identified by three organizing experts, representing realistic and diverse roles, research areas, and backgrounds (e.g., a fifth-year PhD student studying VR currently on the job market; a busy junior professor in NLP who spends much of her time on her startup). For each persona, we specified reply behaviors (responsiveness, email writing style; e.g., slow responder, writes only short emails) and optional issues (potential concerns they might raise with speakers; e.g., requesting additional funding even though their in-person visit is already funded). We split them into two groups and created seminar organizing instances. We set the timeline as a monthly seminar schedule, with the first seminar one month after the simulation start date.
Day 1, Preparation and Setup (30min)
Participants first provided informed consent and completed an AI comfortableness survey assessing their perceptions and expectations with AI automation tools, human coordination tasks, and demographics. They also edited a few seeded templated seminars organizing emails according to their tone to feed the system their writing style. They were then introduced to the DoubleAgents interface and task. One researcher gave a brief walkthrough using a visual diagram and demonstration, covering the system’s core components, including the planning-execution coordination agent loop, respondent agent simulation, policy and context modules, visualization, and the flag mechanism.
Day 1, First System Usage and Evaluation (60min)
Participants then completed a seminar scheduling task (40 minutes), interacting with four simulated speakers with diverse personas, response behaviors, constraints, and coordination tensions. During the task, participants articulated and shared their mental models and usage strategies for DoubleAgents. Afterward, they completed another time of AI comfortableness survey and a system experience survey assessing experience, workload, and attitudes toward specific components such as control, usefulness, and trust. Survey questions were adapted from NASA TLX (Hart, 2006) and prior AI perception, reliance, and long-term evaluation studies (Ross et al., 2023; Basoah et al., 2025; Long et al., 2024). Finally, a short interview asked participants to reflect on their overall experience, including any insights or policies they would like the system to incorporate before Day 2.
Day 2, Second System Usage and Evaluation (60min)
Before users’ next DoubleAgent usages, the researcher first ensured that the system incorporated participants’ first-day planning context and any preferences they had established via policies or system edits. Participants then completed another seminar scheduling task (40 minutes), followed by the same AI perception survey, the system experience survey, and a short interview. Day 2 interviews included follow-up questions comparing participants’ mental models, perceptions, experiences, and expectations of the system between Day 1 and Day 2.
Data Collection and Analysis
We logged participants’ interactions with AI suggestions, including regenerations, rejections, and edits, which were incorporated into DoubleAgents’ implicit learning module and our analysis. Interview data were analyzed using thematic analysis to understand how users perceive the usefulness and trust drive about each feature, and quantitative data were analyzed using paired-samples t-tests to examine whether participants’ experiences and perceptions of the DoubleAgents and AI tools across usage and days.
Appendix C DoubleAgents Prompts
C.1. Coordination Agent
C.1.1. Progress Summary
You are an expert seminar organizer assistant. Your task is to analyze the current progress and return a JSON object. Input information may include: - Speaker details (persona and description) - Possible dates - Scheduling rules - Email log - History of approved plans/actions - Time context - Any user suggestions You should mainly look through Email log and history of approved plans/actions to analyze the current progress. Do not assume any additional actions have been taken. Your JSON output should contain: - summary: A concise summary of the current state, including which days are scheduled, what actions have been taken, and the last date we reached out to each speaker (for tracking and follow-up). - un_or_filled_slots: Two lists of dates — one for filled slots and one for unfilled slots. - pending_speakers: A list of speaker names who have not yet responded or are pending (how many days since we last reached out to them). When mentioning speakers, use the format “English Name (ID)”. Avoid vague terms like *all speakers* or *pending speakers*. ---------- [Task Rules] [Possible Dates] [Time Context] [Speaker Information] Here are the history of Approved Plans and Actions: [Approved Plans and Actions] -------------------------------- Here is the history of past user edits and implicit learning (recent, action edits only): [Implicit Learning] -------------------------------- Past Email Log: [Email Log]
C.1.2. Policy Selection
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy Return a JSON list of the indices of the applicable policies (e.g., [1,2]) or [1]. If none apply, return an empty list []. [Main Reasoning Result] [Coordination Policies] Return ONLY the JSON list of indices. [Implicit Learning/User Edit History (recent, action edits only)]
C.1.3. Action Generation
You are an expert seminar organizer assistant. Your task: Given the organizing progress summary, the selected coordination policies, and the current time context, generate a concise plan for what must be accomplished today only. Output format: Return only one single string that contains: 1. The plan (concise, no bullet points, no unnecessary multiple sentences). 2. A short rationale starting exactly with "Why this plan? ...". The rationale should be about 40 words explaining why this plan was chosen according to the task progress summary context and how the policies/time context informed the decision. Example: "Send email to X to request availability for slots XXX. Why this plan? As we are just beginning the scheduling process and have not reached out to X yet, we followed policy #6 (highlighted on the left) to obtain X’s availability for the slots. This will help jumpstart the process."" Guidelines: - Use each speaker’s real English name. - Focus strictly on today (ignore tomorrow or later). - Valid actions (choose strategically): 1. "Send email to X to request availability for slots XXX" (initial outreach). 2. "Follow up with X" (if email sent but no reply). 3. "Wait for response from X" (if email sent and still within expected waiting window). 4. "Confirm and assign slot for X" (only if X has shared availability; avoid finalizing too early unless time is urgent). 5. "Send clarification email to X" (if clarification on dates/details is needed). - If multiple actions are needed, combine them into one clear sentence (e.g., "Send email to X, Y, and Z to request availability for slots A, B, and C."). - Do not include any other tasks (e.g., slides, addresses, future prep). - Remember what day today is and when the last emailing was done (relevant for some policies). - Consider time urgency: - If few days remain, prioritize immediate action. - If weekend, default to waiting for responses unless a policy specifies otherwise. -------------------------- Time Context: [Time Context] Task Progress Summary: [Progress Summary] -------------------------- Applied Coordination Policies: [Applied Policy] -------------------------- Implicit Learning/User Edit History (recent, action edits only): [Implicit Learning - Action Edits] --------------------------
C.1.4. Email Drafting
I am [Organizer Name], a seminar/workshop organizer. Please help me write an email to [Speaker Name] about [Email Goal]. Please make sure it conveys that. Try to use a casual tone. If I haven’t given you any details, keep it concise and avoid unnecessary formalities. Sometimes it is a full email to invite some, while sometimes it might be just a followup email or a clarification / confirmation email. So you should write different emails according to the situation and past email history. Please return a JSON object with the enhanced email containing: - "from": organzier (just organizer, not names), - "to": recipient (ensure it’s just speaker’s Full name, no other information, dont change) - "subject": subject - "body": enhanced email body that is professional and clear Make sure the email is natural language, professional, and incorporates all context provided. Consider the time urgency when writing the email. If only a few days remain until the seminar dates, convey urgency appropriately. For follow-up emails (like following up / reminder emails -- no need to mention the seminar information.) You need to explicitly mention the dates, don’t say "next seminar/next Thursday" or "Nov 1, 8, 10". You need to list the dates individually. Start the greeting with "Dear [Speaker name]". --------------------------- Here are some information of all the speakers we want to invite Context: [Progress Summary] Past Email History with that Speaker: [Past Email History with Speaker Name] --------------------------- Time Context: [Time Context] --------------------------- Seminar Information (can be opted out if this is not the first email) [Seminar Information] --------------------------- Users’ template examples for invitation, if user has provided some, please try to follow them: User’s history invitation email : [Invitation Email Template] User’s history followup email: [Invitation Followup Template] User’s history confirmation emails : [Invitation Confirmation Template] --------------------------- When necessary, you should adapt some style edit that I have already taken on previous emails. However, be smart on whether they are the correct people to do so. [Implicit Learning - Email Edits]
C.1.5. Implicit User Preference Learning
Plan and Action Edits
You need to describe the user’s edits by comparing the AI/system {item_type} with the user-edited {item_type}, making use of the history context. Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the {item_type} when [context about the stage or progress of the allocation].’ Example: The user removed Tracy and Jack from the AI-generated outreach plan but kept others, such as Tanya and Adam, who are both faculty members. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
Email Edits
You need to describe the user’s email edits related to format and style by comparing the AI/system item_type with the user-edited item_type, making use of the history context. You need to first analyze if the user edits relate to format and style. If not, you should return "No learning needed as the user edits are not related to format and style." Note that you do not need to analyze the user edits if it relates to content (e.g. specific dates) Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the item_type when [context about the stage or progress of the allocation].’ Example: The user changed "Best regards," from the AI-generated outreach plan to "Thanks," in the email when initiating outreach to Jack regarding available dates. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
C.2. Edge Case Detection
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG: SpeakerName seems unavailable due to maternity leave and suggesting a lab substitute. Should I follow up with the proposed colleague or reassign the slot to another invited speaker? Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG: SpeakerName’s traveling during scheduled week requests swap to May. Should I reschedule them to May or retain current slot and reassign to another speaker? Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG: SpeakerName’s uncertain due to family reason, with final notice only one day prior. Should I secure a backup speaker now or wait for their last-minute confirmation? Here are the policies: [Policies]
C.3. Simulated Respondents
You are simulating a seminar speaker. Your persona: [Speaker Persona]. Name: [Speaker Name]. The organizer’s name is: [Organizer Name]. The current simulation day is: [Day ID]. You have received the following email thread (latest at the bottom): [Email Log] Possible dates to choose from: [Speaker Availability]] As the speaker, you can: - Reply to the organizer’s email, providing your availability, asking questions, or giving other relevant information. - If you have questions for the organizer, you can draft them in your reply and wait for a response. - You may choose to wait for more information before replying, if appropriate. - You can reply to multiple emails in the thread if you wish. - You may also choose to wait for a specific number of days before replying (e.g., "wait 2 days and reply on day 5"). If you choose to wait, specify the simulation day you intend to reply. Decide what you want to do next: - If you want to reply, write a detailed, persona-driven reply (possibly including questions for the organizer). - If you want to ask a question and wait for a response, draft your question(s) and indicate you are waiting. Do not ask questions that are too specific (’like equipment setup’ etc, this is a scheduling task, they only care if you are available on the date.) - If you want to wait and not reply yet, explain why, and specify the simulation day you intend to reply (e.g., "wait_until_day": 5). Return your action as one of: ’reply’, ’ask_question’, ’wait’. If you choose to wait, also return a "wait_until_day" field (integer simulation day to reply). Then, provide your reply email (if any) as a JSON object with keys: from, to, subject, body, timestamp. Remember, your subject should be just a short first-point-of-view summary of your reply (less than 10 words, like "I can make it on Sep 5", "I can’t make it on Sep 5", "Sorry I need a few days to think." etc). Also, provide a short "RATIONALE" section (2-4 sentences) explaining your reasoning and thought process for this turn. Format your response as: ACTION: <reply|ask_question|wait> WAIT_UNTIL_DAY: <integer or ’none’> EMAIL: <JSON object or ’none’> RATIONALE: <your reasoning> EMAIL should be like: "from": "XX", "to": "XX", "subject": "XXX", "body": "XXXXX", "timestamp": "XX" dont output json ‘‘‘ etc
C.4. Free-form QA Chat
You are an expert assistant. The following is an email log from a seminar simulation system. Answer the user’s question based on the email log. If the answer is not directly available, say so, but try to be helpful and concise. [EMAIL LOG] [USER QUESTION]
C.5. Daily Summary
Today is simulation day [Day ID]. Here is the email log for today (may include previous days): [Email Log] Please provide a concise 40-word summary of the progress made today, including any key actions taken, responses received, and remaining issues. The summary should be suitable as a notification for users in a sidebar. Dont include today’s date in the summary -- just summarize the progress.
Appendix D Technical Evaluation Details
D.1. Policy Selection
Here are the prompts used in the technical evaluations for policy selection:
D.1.1. Experiment prompts
speaker-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
summary-based
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Main reasoning result: {progress_summary} Coordination Policies: {policy_list} Return ONLY the JSON list of indices:
D.2. Edge Case Detection
D.2.1. Test suite construction
We employ GPT‑5 to generate a set of 10 edge-case (flaggable) test cases and 10 non-flagged (clear) test cases, with all examples validated for quality by the lead authors. To enhance linguistic and contextual diversity, each test case is further expanded into five stylistic variations, reflecting different tones and levels of formality. These variations are conditioned on five distinct speaker personas: a senior renowned faculty member, a junior faculty member, a senior PhD student, a junior PhD student, and a close peer researcher. In total, 100 test cases are constructed and used as our test suite for edge case detection.
D.2.2. Prompts to generate linguistic variations and test data
You are a helpful assistant. You need to generate test cases for me to test my flag detection assistant. I am organizing a seminar series, inviting speakers. You will be acting as the simulated speakers, writing response emails to my invitation email. You will be given cases where the system should mark as clear and that are not edge cases. For each case, you should generate exactly five different variations of responses in the role of the following five speaker personas: 1) a famous distinguished professor who is always busy 2) a junior faculty who is trying to establish herself in the field 3) a senior phd student who is able to graduate and is looking for a job 4) a junior phd student who is just starting his phd journey 5) a close friend of mine who is also a phd student researcher. The following are the cases where the system should mark as clear (i.e. labeled as "0"): Please generate the test cases in the following JSON format: """ { "case_description": $\{"persona":"persona\_description","response":"response","flag":"0$"\} } The case_description is the title of each scenario shown above. Please generate exactly five variations of the response for each case
D.2.3. Experiment prompts
zero-shot
You are an expert seminar organizer assistant. Analyze this response from a speaker invitation. Response: response You need to flag edge-case issues that may require user confirmation or special and urgent attention If an edge-case issue is found, simply output "FLAG", otherwise output "Clear". Your Output:
few_shot without policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Response: {response} Your Output:
few_shot with policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Here are the policies: - "Don’t be shy to ask for more availability if they don’t give you many. They will probably reply more if you beg." - "Some people don’t read or respond to emails right away because they’re busy. You should wait for their response, but it’s okay to follow up every two workday days if you haven’t heard back.", - "If someone hasn’t responded to the email after three reach-outs, then don’t contact them again.", - "Ambiguity is okay. People can typically do the days they are vague about — you can prioritize asking for more availability and just confirm the ambiguity right before assigning them at the end.", - "Some speakers—especially faculty—are more desirable to schedule. We should prioritize their availability by assigning slots to faculty over students and asking students to be more flexible or to provide additional dates to accommodate them.", - "If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots." Response: response Your Output: